### v1

# Configuring Agentic AI Coding Tools: An Exploratory Study

Abstract.
Agentic AI coding tools with autonomous capabilities beyond conversational content generation increasingly automate repetitive and time-consuming software development tasks. Developers can configure these tools through versioned repository-level artifacts such as Markdown and JSON files. In this paper, we present a systematic analysis of configuration mechanisms for agentic AI coding tools, covering Claude Code, GitHub Copilot, Cursor, Gemini, and Codex. We identify eight configuration mechanisms and, in an empirical study of 2,926 GitHub repositories, examine whether and how they are adopted. We then explore Context Files, Skills, and Subagents, that is, three mechanisms available across tools, in more detail. Our findings reveal three trends. First, Context Files dominate the configuration landscape and are often the sole mechanism in a repository, with AGENTS.md emerging as an interoperable standard across tools. Second, advanced mechanisms such as Skills and Subagents are only shallowly adopted: most repositories define only one or two artifacts, and Skills predominantly rely on static instructions rather than executable workflows. Third, distinct configuration cultures are forming around different tools, with Claude Code users employing the broadest range of mechanisms. These findings establish an empirical baseline for longitudinal and experimental research on how configuration strategies evolve and affect agent performance as agentic AI coding tools mature.
Software Engineering, Generative AI, AI Agents, Configuration
journalyear: 2026copyright: ccdoi: XXXXXXX.XXXXXXXdoi: XXXXXXX.XXXXXXXconference: Anonymous Conference; Anoym. Conf.; Earthbooktitle: Anonymous Conferencedoi: 10.1145/XXXXXXX.XXXXXXXisbn: 979-8-4007-XXXX-X/2026/04ccs: Software and its engineering
1. Introduction
Agentic AI coding tools based on large language models (LLMs) (Sergeyuk et al., 2025; Terragni et al., 2025), automate time-consuming and repetitive tasks, such as generating and editing code, tests, and documentation. The agentic capabilities of these tools go beyond those of purely reactive conversational assistants. Instead of just following user commands, these tools take initiative and act proactively to accomplish defined objectives (Anthropic, 2025). They can autonomously interact with development environments and use project artifacts, external data, and command-line tools to perform actions with reduced human interaction (Yang et al., 2024). Certain functionality can be delegated to “tools” and “(sub)agents.” Tools are capabilities with a specific bounded function (e.g., run_python(code) or search_web(query)). [^1] We distinguish such “tools” from full-fledged “agentic tools” (e.g., Claude Code). They are deterministic and invoked by the underlying model or agents. (Sub)agents are goal-directed action-taking components. They interpret a user goal, break it down into substeps, choose appropriate tools, execute actions, observe results, and potentially adjust their plan for the next iteration of this “agent loop.” Initially, agentic AI coding tools implemented one central agent loop steered by a foundational model such as Anthropic’s Opus or OpenAI’ GPT. Example tools include Claude Code and OpenAI Codex. Soon, conversational tools such as GitHub Copilot and Cursor offered similar capabilities in an “agent mode.” Over the months preceding the writing of this paper, tool vendors implemented extension and configuration mechanisms that allow developers, for example, to write their own (sub)agents that operate in parallel to the central agent loop.
We define configuration mechanisms as means by which developers can tailor tool and agent behavior to a project or workflow (e.g., by providing contextual information in a file or by specifying dedicated subagents). A configuration artifact is the tangible specification of that mechanism (e.g., the actual Markdown file with the contextual information). For example, files like AGENTS.md or CLAUDE.md act as “READMEs for agents” with context-specific information about build commands, coding conventions, and rules for CI/CD pipelines (Mohsenimofidi et al., 2026; OpenAI, 2025). Configuration artifacts are version-controlled, inspectable, and collaboratively maintained. With the increased availability and diversity of agentic tools, the number of configuration mechanisms and related artifacts has also increased.
To understand the current landscape of agentic tools’ configuration mechanisms and their use in open-source software (OSS), we address the following research questions:
RQ1:
What configuration mechanisms do agentic coding tools offer?
RQ2:
Which mechanisms are adopted in OSS repositories?
RQ3:
How are configuration mechanisms adopted?
We consider configuration mechanisms that are captured in configuration artifacts to be consumed by agentic tools, are repository-versioned for collaborative maintenance, and include instructions intended to customize the behavior of AI coding tools.
With this paper, we provide the following contributions: (1) We systematically documented eight configuration mechanisms: Context Files, Skills, Subagents, Commands, Rules, Settings, Hooks, and MCP servers. We identified these from the documentation of Claude Code, GitHub Copilot, Cursor CLI, Gemini CLI and Codex CLI. Some mechanisms are available in all tools; others are specific to particular tools. (2) We analyzed the adoption of these configuration mechanisms in 2,926 GitHub repositories. Context Files (Markdown files that provide contextual project information) dominated and are often the sole configuration mechanism. Claude Code users apply the broadest range of configuration mechanisms. (3) We analyzed the adoption of Context Files, Skills (bundles of information and tools), and Subagents (specialized agents with their own model context), in more detail. Three Context File formats showed the most dynamic development, with AGENTS.md emerging as a standard. Although Skills and Subagents can be defined for a wide range of purposes, most repositories that adopt them define only one or two artifacts. Moreover, Skills primarily rely on static resources rather than executable scripts to extend tool and agent behaviour.
Figure 1. Data collection process.
Diagram showing the data collection process.
2. Related Work
Related work has examined how repository-level context files, prompts, and structured context mechanisms shape agent behavior, efficiency, and integration into development workflows. Although several studies have investigated individual configuration artifacts, most notably repository-level context files such as AGENTS.md, prior work has typically focused on single artifact types in isolation. There has been no systematic examination of the broader configuration mechanisms to customize agentic behavior. Independent of tool architecture, context engineering allows developers to design, structure, and supply task- and project-specific information to guide the behavior of tools (Mei et al., 2025; Schmid, 2025; Horthy, 2025; Sahoo et al., 2024; DAIR.AI Prompt Engineering Guide, 2025). Empirical work has mainly focused on repository-maintained context files, such as AGENTS.md. Mohsenimofidi et al. (Mohsenimofidi et al., 2026) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level instruction files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions. Complementing this work, Chatlatanagulchai et al. (Chatlatanagulchai et al., 2025) found that context files are actively maintained, structurally consistent, and focus predominantly on functional development instructions such as build and testing procedures. Santos et al. (Santos et al., 2025) analyzed Claude Code projects and how configuration artifacts structure agent interactions in repositories.
Beyond structural characterization, recent work has begun to evaluate the impact of such artifacts. Lulla et al. (Lulla et al., 2026) conducted a controlled study comparing agent executions with and without an AGENTS.md file and reported lower runtime and token consumption while maintaining comparable task completion behavior.
In other work, Villamizar et al. (Villamizar et al., 2025) argued that prompts should be treated as software engineering artifacts and outlined a research agenda to understand their evolution, reuse, and governance. Beyond static prompts, Zhang et al. (Zhang et al., 2025) introduced a framework that treats contexts as evolving instructions and addresses issues such as brevity bias and context collapse through incremental updates.
Recent work further generalized this perspective by formalizing context as a programmable abstraction. Zhang and Wang (Zhang and Wang, 2025) introduced Monadic Context Engineering, modeling context construction and transformation through monadic composition. McMillan (McMillan, 2026) evaluated structured file-native context schemas and showed how format and multi-file organization influence retrieval accuracy and downstream agent performance. Ye et al. (Ye et al., 2026) proposed Meta Context Engineering, a framework in which agentic skills and context artifacts co-evolve via evolutionary search.
However, a cross-tool and cross-agent perspective mapping configuration mechanisms and studying their adoption patterns across repositories is missing. Our work addresses this gap.
3. Data Collection and Analysis
We collected OSS projects hosted on GitHub. Since GitHub hosts a variety of content, we selected repositories belonging to “engineered” software projects (Munaiah et al., 2017). We adopted and updated a selection approach used in prior work, starting with the SEART GitHub search tool  (Mohsenimofidi et al., 2026; SEART, 2025; Dabic et al., 2021). We selected non-fork repositories with at least two contributors and a license that were created before 1 January 2024 with commits since 1 June 2025. We then excluded archived, disabled, or locked repositories We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and applied filters based on commit and watcher count based on prior research (Mohsenimofidi et al., 2026). This resulted in a final sample of 37,249 repositories.
Table 1. Overview of repository-level configuration mechanisms across agentic AI coding tools. Each cell lists the repository file(s) or directory implementing that mechanism; “–” indicates that the mechanism is not available.
Mechanism Description Claude Copilot Codex Cursor Gemini
Context Files Markdown file providing persistent context loaded every session. CLAUDE.md .github/ {
copilot-instructions.md
| instructions/*.md }a
AGENTS.md,
AGENTS.override.md
AGENTS.md, .cursorrulesc GEMINI.md
Settings JSON/TOML config for project-level tool behavior. .claude/ settings (local)?.json b .codex/ config.toml .cursor/ cli.json .gemini/{ settings.json | config.yaml }
Skills Reusable knowledge and invocable workflows. .claude/skills/ .github/skills/ .codex/skills/ .cursor/skills/ .gemini/skills/
Subagents Specialized agents running in an isolated context. .claude/agents/ .github/agents/ .cursor/agents/
Commands User-triggered shortcuts for predefined prompts. .claude/ commands/ .cursor/ commands/ .gemini/ commands/
Hooks Scripts executed at specific agent lifecycle points. .claude/ settings.json .github/hooks/*.json .cursor/ hooks.json .gemini/ settings.json
Rules System-level instructions to control agent behavior. .codex/rules/ .cursor/rules/
MCP External tool or data connections via the Model Context Protocol. .mcp.json b .codex/ config.toml .cursor/ mcp.json .gemini/ settings.json
a Copilot also supports CLAUDE.md, AGENTS.md, and GEMINI.md; b Configured via the GitHub web interface, not via files in the project repository;
c Cursor deprecated .cursorrules and now suggests using AGENTS.md instead.
Table 2. Release dates of agentic AI coding tools and repositories ($n=2,926$; one repository can use multiple tools).
Agentic Tool Release (Month/Year) #Repositories
Claude Code 02/2025 (CLI & agents since release) 1,310
GitHub Copilot 10/2021 (Release)
02/2025 (Copilot Agent Mode)
09/2025 (Copilot CLI)
958
Codex CLI 04/2025 (CLI & agents since release) 558
Cursor CLI 03/2023 (Release)
06/2025 (Cursor Agents)
08/2025 (Cursor CLI)
355
Gemini CLI 02/2024 (Release)
05/2025 (Gemini Agent Mode)
06/2025 (Gemini CLI)
186
Table 3. Repository metadata comparison by agentic tool. Cells show median, IQR, and Cliff’s $\delta$ for tool-adopting repositories. Significant differences (BH-adjusted) are in bold.
Agentic tool Age (years) Contrib. Commits Size (KB)
All 6.7 42 2,106 40k
(nn=2,926) (4.3–9.4) (19–105) (965–5,064) (9,498–132k)
Claude 6.2∗∗∗ 40 2,234 42k
(nn=1,310) (4.1–9.2) (19–111) (995–5,268) (10k–151k)
$\delta$=-0.08 $\delta$=0.00 $\delta$=0.03 $\delta$=0.04
Gemini 6.8 66∗∗∗ 3,322∗∗∗ 58k∗∗
(nn=186) (4.7–9.5) (32–144) (1,290–9,776) (15k–193k)
$\delta$=0.03 $\delta$=0.21 $\delta$=0.21 $\delta$=0.12
Codex 6.9 43 1,809∗∗ 34k∗∗
(nn=568) (4.2–9.3) (19–89) (877–4,058) (8,696–94k)
$\delta$=0.01 $\delta$=-0.02 $\delta$=-0.10 $\delta$=-0.08
Copilot 7.1∗∗∗ 42 2,200 44k
(nn=958) (5.0–9.7) (19–113) (1,000–5,591) (9,858–143k)
$\delta$=0.11 $\delta$=0.02 $\delta$=0.04 $\delta$=0.03
Cursor 5.5∗∗∗ 54∗∗ 2,780∗∗∗ 75k∗∗∗
(nn=355) (3.4–7.9) (23–120) (1,264–5,968) (21k–201k)
$\delta$=-0.19 $\delta$=0.10 $\delta$=0.14 $\delta$=0.22
Note: Mann–Whitney U test (two-sided) with Benjamini–Hochberg FDR correction (20 comparisons). Effect sizes: Cliff’s delta. $p<.05$; ∗∗ $p<.01$; ∗∗∗ $p<.001$ (adjusted).
Figure 2. Adoption of agentic tools per programming language; % relative to #repositories using a certain tool.
Figure 3. Usage of configuration mechanisms across agentic tools. % relative to #repositories using a certain mechanism.
Diagram showing adoption of AI coding tools per programming language.
Figure 1 outlines our data collection pipeline for these repositories. We cloned them and searched their default branch for a README file. We excluded 11 repositories that became unavailable in the meantime, 60 repositories without a README file, and then used the lingua-language-detector Python library to detect the language of files we found. This resulted in 994 repositories with non-English README files being excluded.
For the remaining 36,184 repositories, we developed a classification pipeline that uses the GPT-5.2 model to determine, based on the content of a README file, whether the software project fulfills our definition of an “engineered” project. We based our definition on Munaiah et al.’s paper (Munaiah et al., 2017). An engineered project is a project that shows clear evidence of software engineering practices (e.g., it includes a clearly stated purpose of a problem that the software solves; it provides installation and deployment instructions, developer and user documentation, testing, CI/CD, or quality-assurance-related information). There should also be evidence for maintenance. This definition was part of our prompt, together with the README of each repository. In our supplementary material, we share the final prompt and previous iterations of the prompt, which we tested with subsets of the data, as well as the exact configuration for labeling repositories as “engineered” or not (we kept OpenAI’s default values for all parameters) (Authors, 2026). The classification pipeline labeled 32,564 repositories as engineered software projects, was ‘unsure’ about 2,204 repositories, and excluded 1,416 repositories. For each label category, we spot-checked randomly selected samples, both during the iterative development of the prompts and for the final sample.
We then cloned the remaining 32,564 repositories and applied heuristics based on file names and file paths to detect the usage of AI coding tools and configuration mechanisms (see Table 1). These heuristics are based on the answers to RQ1 and are briefly discussed in Section 4. This resulted in 2,926 repositories that use one or more AI coding tools. Our data collection and analysis scripts and the analyzed data are available online (Authors, 2026).
4. Configuration Mechanisms (RQ1)
To determine the set of configuration mechanisms for agentic AI coding tools, we followed two steps. We first selected five agentic AI coding tools based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026). The tools represent the four most popular AI tools among the surveyed developers. We added Codex instead of ChatGPT as Codex is an agentic tool by the same vendor using the same models. We added Cursor because it has recently been studied in software engineering  (He et al., 2026; Jiang and Nam, 2026). Table 2 lists the five tools. Our study focuses on agentic tools with a command-line interface (CLI) that first appeared in 2025. For some tools, prior non-agentic versions exist.
In the second step, one author systematically reviewed the online documentation of each tool, documenting configuration mechanisms together with repository-level files and directories that might indicate their usage. This was then cross-checked by two other authors and discussed among all.
As part of RQ1 and to “operationalize” configuration mechanisms and their occurrence in repositories, we then used that information to develop heuristics to identify whether tools are used in a repository (see Table 1). Note that for GitHub Copilot, Cursor, and Gemini, files and directories indicating their use apply both to their conversational and agentic interfaces. For example, a .claude folder or a CLAUDE.md file indicate usage of Claude Code. In addition, we developed strategies to determine whether configuration mechanisms were used. For example, if a directory .claude/agents/ with Markdown files is present, this indicates usage of Subagents. All documents and the Python scripts that implement our matching strategy are part of the supplementary material (Authors, 2026).
{results}
Key insights for RQ1:
  • Qualitative analysis of the documentation of agentic tools revealed eight configuration mechanisms.
  • There is convergence for some mechanisms across tools (e.g., all five support Context Files and Skills), yet no single tool implements all eight mechanisms, suggesting the ecosystem is still maturing, standards are developing.
5. Adoption of Configuration Mechanisms (RQ2)
To understand how developers adopt configuration mechanisms in agentic AI coding tools, we analyze the presence and co-occurrence of configuration artifacts across repositories.
5.1. Characterization of Repositories in Dataset
Before analyzing adoption, we characterize the 2,926 repositories in our dataset to contextualize the configuration mechanisms of agentic AI coding tools.
Most repositories adopted a single tool. Among repositories with multiple tools, Claude appeared most frequently with others. The most common combination was Claude and Copilot ($n=128$). Overlaps of three or more tools were rare. In our initial sample, the top five main programming languages per repository ($n=36,194$) were Python (8,133; 22.5%); TypeScript (4,999; 13.8%); Java (3,813; 10.5%); Go (3,786; 10.5%); and JavaScript (3,709; 10.3%). Interestingly, this order is slightly different for repositories that use agentic coding tools ($n=2,926$, see Figure 3) where TypeScript and Go were more prominent and C# replaces JavaScript: TypeScript (922, 31.5%); Python (548, 18.7%); Go (533, 18.2%); C# (256, 8.8%); Java (246, 8.4%). TypeScript is the most common primary language for repositories using Claude, Codex, Copilot, and Cursor. Usage in Java and C# repositories is lower for all tools but Copilot. As shown in Table 3, repositories using Cursor are younger compared to those using other tools. Repositories associated with Gemini exhibit larger contributor counts and commit volumes, compared to those associated with other tools. Cursor repositories are the largest in terms of kilobytes of source code.
5.2. Distribution of Configuration Mechanisms
Figure 3 shows the distribution of configuration mechanisms. Across repositories, Context Files (with artifacts such as CLAUDE.md or AGENTS.md files) were the most frequently adopted configuration mechanism. In contrast, other configuration mechanisms, including Rules, Settings, Commands, Skills, and Subagents, were substantially less prevalent. Each of these mechanisms appeared in fewer than 20% of the repositories. Rules are primarily concentrated in repositories using Cursor, since Cursor was one of the first to introduce Rules (Jiang and Nam, 2026).
Most repositories adopt a limited subset of available configuration mechanisms. In many cases, repositories include only a single Context File artifact. The adoption of multiple non-context file configuration mechanisms within the same repository remains comparatively rare. These results suggest that configuration usage is currently characterized by one dominant baseline configuration (i.e., Context Files).
Figure 4. Cumulative adoption of selected configuration artifacts for agentic tools. Copilot Instructions include the two different artifact types for Context Files, see Table 1.
Figure 5. Configuration mechanism count per repository.
Diagram showing the comulative adoption of configuraiton artifacts for agentic tools.
5.2.1. Co-occurrence of Configuration Mechanisms
Several configuration mechanisms are frequently adopted together. Some of these correlations can be explained by shared configuration artifacts. For example, Settings and Hooks have a strong positive correlation ($\rho=0.52$), because hooks are usually defined in Settings files. However, Settings and Hooks also both co-occur with Skills ($\rho=0.23$ and $\rho=0.24$, respectively), which are defined and configured differently. Subagents also show positive correlations with multiple mechanisms, including Settings, Skills, and MCP. Other configuration mechanisms appear less frequently in combination. Context Files are negatively correlated with Rules ($\rho=-0.36$) and Commands ($\rho=-0.14$), which can again attributed to how they are defined and which tools first introduced the mechanisms. Overall, configuration mechanism adoption across repositories is characterized by recurring combinations of mechanisms rather than uniform usage. Patters of co-occurrence might be impacted by the tools and configuration mechanisms they support (e.g., Claude, the dominating tool, does not support Rules).
5.2.2. Adoption of Configuration Mechanisms over Time
In Figure 5, we show the adoption of configuration mechanisms over time.
The figure shows individual artifacts that together form the mechanism Context Files, plus Skills and Subagents which have a uniform format.
Context Files (CLAUDE.md, AGENTS.md, copilot-instructions.md) clearly dominate and increase continuously, while Skills and Subagents experienced a comparatively slow growth. As the figure shows, .cursorrules and copilot-instructions.md started being introduced in 2024. Although Cursor and also Copilot were originally released in 2023 and 2021, respectively, agentic capabilities were only introduced in 2025 (see Table 2). The adoption of copilot-insturctions.md has increased since then.
{results}
Key insights for RQ2:
  • Context Files are the dominant and often sole configuration mechanism. Advanced configuration mechanisms (e.g., Hooks, MCP) are adopted less frequently.
  • Adoption varies: Claude has a broad configuration footprint, while Cursor emphasizes rule-based mechanisms.
  • Although certain correlations between configuration mechanisms exist, the landscape still appears in flux.
6. Details of Configuration Mechanisms (RQ3)
In this section, we analyze the use of three configuration mechanisms in more detail. We selected Context Files and Skills because they are supported by all tools, as discussed in Section 5, and Subagents because they follow a similar format as Skills.
6.1. Configuration Mechanism: Context Files
Context Files are Markdown files that provide a central machine-readable source of contextual information about the repository in which they are used. AGENTS.md was introduced by OpenAI and now serves as an open tool-agnostic convention that is increasingly supported by various tools (Mohsenimofidi et al., 2026). We analyzed the presence and co-occurrence of Context Files across repositories. We identified 4,860 Context Files across 2,634 of the 2,926 repositories in our sample. The histogram in Figure 5 shows the distribution of Context Files. Most repositories that do use Context Files include one or two of such files.
CLAUDE.md emerges as the dominant file type with 1,661 (34.2%), followed closely by AGENTS.md and copilot-instructions.md with 1,572 (32.3%) and 1,344 (27.7%) files, respectively. GEMINI.md (159 files, 3.3%) and .cursorrules (73 files, 1.5%) are rare. Note that .cursorrules are now deprecated; Cursor suggests using AGENTS.md instead. Cursor rules are different and are stored in .cursor/rules/. We also investigated the repository-level adoption of the Context Files, where CLAUDE.md demonstrates the highest adoption rate with 45.40% of repositories (1,195 repos), AGENTS.md follows closely with 40.6% adoption (1,069 repos), and copilot-instructions.md appears in 35.1% of repositories (925 repos). The predominance of CLAUDE.md, AGENTS.md and copilot-instructions.md, both in terms of their frequency and the number of repositories adopting them, indicates a clear concentration of projects around these files.
We then investigated programming language-specific patterns in Context Files usage across the repositories. We found Context Files in 651 repositories with TypeScript as the main language, 435 with Python, 414 with Go, 210 with Java, 210 with C#, 201 with JavaScript, 134 with Ruby, 181 with Rust, 156 with C++, 123 with PHP, and 53 with C. CLAUDE.md is the most common Context File across languages, except for Java, C#, and C++, where copilot-instructions.md emerges as the dominant type.
Figure 6. Creation order of Context Files per repository. Curly braces indicate that files were added on the same day.
Figure 7. Reference-only Context Files. Examples of other files include README.md and CONTRIBUTING.md.
Diagram showing the creation order of Context Files per repository.
Figure 7 shows the creation order of the different Context Files. Most repositories only used one type of Context Files. These repositories are not captured in this figure. AGENTS.md and CLAUDE.md are typically created first, and if CLAUDE.md existed first, then AGENTS.md was commonly created afterward. A reason for this could be that Claude Code is the most popular agentic AI coding tool, but does not yet support the emerging standard AGENTS.md (anthropics/claude-code on GitHub, 2026). Interestingly, repositories that started with copilot-instructions.md then often created CLAUDE.md or AGENTS.md files, even though GitHub Copilot supports all major Context Files types. This, in addition to the overall dominance of CLAUDE.md and AGENTS.md, illustrates that these two files currently serve as the de facto standard.
Finally, we observed that Context Files commonly refer to each other. This is visualized in Figure 7. “Referring to” means that one file, the source, contains a reference to another file, the target, rather than providing their own content. We found three types of references between Context Files:
  1. (1)
    Direct pointer: Context Files include one line that references files either via their filename (e.g., AGENTS.md, @./AGENTS.md), imperative statements (e.g., Read @AGENTS.md), or a Markdown link (e.g., [anchor text](AGENTS.md)).
  2. (2)
    Short reference with context: Context Files provide 2–5 lines of what to do with the referenced file, e.g., Use instructions from AGENTS.md to guide your work., Always read AGENTS.md before answering, or a Markdown header followed by one of the above.
  3. (3)
    Brief summary plus reference: Context Files include a header, 1–2 sentences of context, and a reference to another file, e.g., # AI Guidelines […] **Read @AGENTS.md for comprehensive guidelines.**
We found 518 file pairs that reference each other. The file type with most references was CLAUDE.md, with 357 “outgoing” references, followed by AGENTS.md and GEMINI.md. In contrast, the most frequently referenced file was AGENTS.md with 368 “incoming” references, substantially more than CLAUDE.md and copilot-instructions.md. Furthermore, the strongest pattern was the CLAUDE.md to AGENTS.md pair that occurred 311 times. In contrast, when AGENTS.md files acted as a pointer, they most frequently referenced CLAUDE.md files. This again reflects the symbiotic relationship between these two types of Context Files that we also found in the creation order discussed earlier.
6.2. Configuration Mechanism: Skills
Skills were introduced by Anthropic and now serve as an open standard for extending the capabilities of agentic AI tools with specialized contextual knowledge and workflows. Skills bundle prompts, tools, and documentation that an agent can invoke on demand. From an implementation perspective, Skills are a directory that contains a SKILL.md file. This file must include a YAML frontmatter with at least a name and description, followed by instructions in the body for how a task should be performed.
We found 601 Skills in 158 repositories. On average, these repositories use 3.8 Skills ($min=1.0$; $max=28.0$; $median=2$). However, as Figure 5 shows, repositories that use Skills typically use fewer than three Skills.
As with Context Files, we also explored whether there are programming language-specific patterns. Figure 3 shows that TypeScript and Python are the dominating programming languages in all repositories that use agentic tools. This distribution is also reflected across projects that use Skills.
The specification of Skills recommends that SKILL.md files should contain fewer than 500 lines. It further recommends moving any detailed material to separate files (agentskills.io, 2026). We found that of the 601 Skills we identified, only 25 (4%) were longer than 500 lines. That is, most developers follow the recommendation. To provide detailed information and instructions, Skills can utilize different resources, i.e., additional optional files or directories that are loaded by the agents of a tool when required. There are three types of resource:
  1. (1)
    scripts/ contains executable code that agents can run. Language options include Python, Bash, and JavaScript.
  2. (2)
    references/ contains documentation that an agent can read when needed. This can include technical references, templates or structured data and other domain-specific files.
  3. (3)
    assets/ contains static resources, such as templates (document templates, configuration templates), images (diagrams, examples) and data files (lookup tables, schemas).
To obtain the resources used by the different Skills, we scanned all Skills directories in the studied repositories for the occurrence of the relevant resource folders. Of the 601 skills analyzed, the vast majority (501, 83.3%) included no additional resources. Among those that did, the most common pattern was a references/ directory (48 skills, 8.0%), followed by a scripts/ directory (34, 5.7%). Twelve Skills (2.0%) combined both scripts/ and references/, while an assets/ directory appeared in only four Skills (0.7%). The remaining combinations, i.e., references/ with assets/, and all three directories together, each occurred just once (0.2%). In summary, Skills mostly use “static” resources (documentation that the agent reads when needed) rather than “dynamic” resources, such as executable scripts that extend agent behavior.
6.3. Configuration Mechanism: Subagents
Subagents are specialized AI agents that agentic AI coding tools or other agents can delegate tasks to. Technically, Subagents are a Markdown file with the same YAML frontmatter structure as Skills. However, Subagents differ in that they operate within their own context window and return the results to the parent agent whereas Skills execute instructions within the calling agent’s context (Anthropic, 2026; Cursor, 2026).
We found 452 Subagents across 131 repositories. On average, these repositories use 3.45 Subagents ($min=1.0$; $max=18.0$; $median=2$). However, Figure 5 indicates that these repositories commonly use fewer than three Subagents. One feature of Claude Code’s Subagents is that they can have their own “memory”, i.e., a persistent directory that survives across interactions with the agent (Anthropic, 2026). This directory can be used to build up knowledge over time, e.g., to store debugging insights. However, we could not find any repositories storing such memory files.
{results}
Key insights for RQ3:
  • Context Files: CLAUDE.md, AGENTS.md and copilot-instructions.md are the most common artifacts. Creation and reference patterns revealed that CLAUDE.md and AGENTS.md serve as de facto standards while AGENTS.md emerges as a unifying standard.
  • Skills: Although Skills can be fine-grained and offer various customization features, most repositories that adopt them only define one or two and predominantly utilize static documentation rather than dynamic scripts.
  • Subagents: The usage patterns for Subagents are similar to those of Skills. Future work should study differences in content and use cases in more detail.
6.4. Threats to Validity
We organize this section following established guidelines (Runeson and Höst, 2009).
Construct validity:
Our heuristics (Table 1) detect the presence of configuration artifacts but not whether the corresponding tool is actively used. We partially mitigate this by restricting our sample to actively maintained repositories, but artifact presence remains a proxy for adoption. Additionally, for GitHub Copilot, Cursor, and Gemini, detected files apply to both conversational and agentic modes, so we cannot isolate agentic usage. This limitation does not affect agentic-only mechanisms such as Skills and Subagents, nor the broader AGENTS.md trend.
Internal validity:
We classified repositories as “engineered” software projects using GPT-5.2 on their README content, with iterative prompt development and spot-checks to mitigate errors. We used a single labeling run without considering inter-model agreement; 2,204 “unsure” cases (often due to inaccessible linked resources) were excluded. Future work should assess the reliability of this approach by adding alternative models and configurations. Detection heuristics were designed by one author and cross-checked by two others, but rapidly evolving tool ecosystems may introduce conventions not yet captured.
External validity:
Our study covers only open-source repositories on GitHub; practices in proprietary or enterprise settings may differ. We selected repositories showing established engineering practices, but cannot claim representativeness for closed-source development, nor did we examine variation across application domains. Finally, our findings are a point-in-time snapshot (February 2026) of a rapidly evolving landscape.
7. Discussion
Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,926 repositories. Below, we discuss the implications of our study.
Standardization around AGENTS.md:
Our results reveal a clear trajectory toward AGENTS.md as a unifying, tool-agnostic configuration artifact. The creation order and reference patterns point in the same direction: For example, when both files are present, CLAUDE.md typically appears first and AGENTS.md is added later (Figure 7). CLAUDE.md most frequently points to AGENTS.md (311 cases), and AGENTS.md receives 368 incoming references overall, far more than any other file type. Repositories that began with copilot-instructions.md often later introduced CLAUDE.md or AGENTS.md, despite GitHub Copilot already supporting all major Context File types. Together, these patterns suggest bottom-up convergence around AGENTS.md as a cross-tool baseline, driven by developer practice and compatibility needs rather than vendor mandate. At the same time, the layering of multiple context files within a single repository raises the risk of redundant or conflicting instructions across artifacts. For tool vendors, native support for AGENTS.md—as already provided by Cursor and Codex—may become a baseline expectation.
Shallow adoption despite available depth:
Across all three configuration mechanisms we examined in detail, adoption remains limited. Most repositories include only one or two Context Files (Figure 5), and the median number of Skills and Subagents per repository is two in both cases. Moreover, 83.3% of Skills do not include additional resources, and when resources are present, static documentation (references/) is more common than executable scripts (scripts/). In practice, Skills, therefore, function primarily as structured text rather than executable workflow bundles. This suggests that configuration is currently used more as documentation than as automation. We also found no evidence of repositories using the persistent memory feature of Claude Code’s Subagents.
This gap between available depth and actual usage likely reflects both the novelty of these mechanisms and the effort required to configure them. Claude Code was released only in February 2025, with Skills and Subagents introduced even more recently. Beyond recency, developers may gravitate toward the lowest-friction mechanism (i.e., Context Files) without exploring more advanced options. Defining executable Skills with scripts and structured resources requires additional design and maintenance effort compared to authoring Markdown instructions, which may deter adoption in the absence of clear guidance on best practices. At present, there is little empirical evidence on which configuration strategies are most effective or under which conditions they yield measurable improvements. Future work should therefore assess whether deeper configuration leads to performance gains, extending early evidence on the impact of Context Files (Lulla et al., 2026).
Distinct tool ecosystems:
Configuration practices vary systematically across tools. Claude Code repositories exhibit the broadest configuration footprint, whereas Cursor projects emphasize Rules and Commands, and Copilot and Codex repositories rarely extend beyond Context Files. These differences likely reflect both the configuration options each tool exposes (Table 1) and the norms emerging within their respective user communities. Tool ecosystems also differ in repository characteristics: Cursor repositories tend to be younger and larger, while Gemini repositories show comparatively higher activity levels (Table 3). An open question is whether these tool-specific configuration cultures will converge as feature sets overlap, or whether distinct patterns of use will persist.
Practical implications and future directions:
Our systematically compiled overview of eight configuration mechanisms (Table 1) provides a reference for developers seeking to understand the options available to them. Based on our findings, we offer the following concrete observations for practitioners. First, Context Files—particularly AGENTS.md—represent the lowest-friction entry point for configuring agentic tools and are already widely adopted. Second, developers who rely on multiple tools should consider maintaining an AGENTS.md file as a shared configuration baseline, given its cross-tool support and the reference patterns we observed. When multiple context files coexist, teams may benefit from clearly structuring them hierarchically (e.g., using tool-specific files as adapters that reference a shared core file) to reflect the layering patterns we identified. Third, for recurring, well-defined workflows, Skills offer the potential for richer configuration through scripts and structured resources, but this potential is currently underutilized. In practice, configuration is presently used more as documentation than as automation, suggesting that teams should adopt executable Skills deliberately and only when the expected benefits justify the additional setup effort. Finally, given the distinct configuration profiles across tools, teams should be mindful that adopting tool-specific mechanisms (e.g., rule-based configurations) may shape their workflows in tool-dependent ways.
For researchers, our findings point to several directions. Longitudinal studies are needed to track how configuration practices evolve as tools mature and as developers gain experience. Controlled studies should assess whether advanced mechanisms such as Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone, extending prior work on the operational impact of AGENTS.md (Lulla et al., 2026). Given that projects increasingly use multiple tools (Section 5.1), research should also investigate interoperability challenges: how configuration artifacts interact across tools, and how potential conflicts between overlapping or contradictory instructions in different files can be detected and resolved. Our automated data collection and analysis pipeline (Authors, 2026) supports such ongoing analyses.
Finally, our study is limited to a single point-in-time snapshot of a rapidly evolving landscape. Agentic AI coding tools are under active development, and both the available configuration mechanisms and the patterns of their adoption are likely to change. The trends we identify—toward standardization around AGENTS.md, shallow adoption of advanced mechanisms, and tool-specific configuration cultures—are early empirical signals rather than settled findings.
8. Conclusions
Our study is the first to provide a comprehensive overview of the configuration mechanisms used by agentic AI coding tools in software engineering. We systematically identified eight configuration mechanisms across five agentic AI coding tools and analyzed their adoption in 2,926 GitHub repositories. Three findings stand out. First, Context Files dominate and are often the only mechanism present, with AGENTS.md emerging as an interoperable standard. Second, advanced mechanisms such as Skills and Subagents remain shallowly adopted, with Skills used predominantly as static documents rather than executable workflows. Third, distinct configuration cultures are forming around different tools. For practitioners, these findings suggest that AGENTS.md is the natural starting point for configuring agentic tools, especially in multi-tool environments, and that Skills offer untapped potential for encoding recurring workflows beyond static instructions. Tool providers should consider improving onboarding and documentation for advanced mechanisms, given the gap between their expressive power and current adoption. For researchers, controlled studies are needed to determine whether richer configuration, such as Skills with executable resources, yields measurable improvements over Context Files alone, and how conflicts between configuration artifacts in multi-tool repositories can be detected and resolved. Longitudinal research needs to track how configuration patterns evolve.

### v2

@@
 Abstract.
-Agentic AI coding tools with autonomous capabilities beyond conversational content generation increasingly automate repetitive and time-consuming software development tasks.
+Agentic AI coding tools increasingly automate software development tasks.
 Developers can configure these tools through versioned repository-level artifacts such as Markdown and JSON files.
-In this paper, we present a systematic analysis of configuration mechanisms for agentic AI coding tools, covering Claude Code, GitHub Copilot, Cursor, Gemini, and Codex.
-We identify eight configuration mechanisms and, in an empirical study of 2,926 GitHub repositories, examine whether and how they are adopted.
-We then explore Context Files, Skills, and Subagents, that is, three mechanisms available across tools, in more detail.
-Our findings reveal three trends.
+We present a systematic analysis of configuration mechanisms for agentic AI coding tools, covering Claude Code, GitHub Copilot, Cursor, Gemini, and Codex.
+We identify eight configuration mechanisms spanning a spectrum from static context to executable and external integrations, and, in an empirical study of 2,923 GitHub repositories, examine whether and how they are adopted, with a detailed analysis of Context Files, Skills, and Subagents.
 First, Context Files dominate the configuration landscape and are often the sole mechanism in a repository, with AGENTS.md emerging as an interoperable standard across tools.
-Second, advanced mechanisms such as Skills and Subagents are only shallowly adopted: most repositories define only one or two artifacts, and Skills predominantly rely on static instructions rather than executable workflows.
+Second, advanced mechanisms such as Skills and Subagents are only shallowly adopted.
+Most repositories define only one or two artifacts, and Skills predominantly rely on static instructions rather than executable workflows.
 Third, distinct configuration cultures are forming around different tools, with Claude Code users employing the broadest range of mechanisms.
-These findings establish an empirical baseline for longitudinal and experimental research on how configuration strategies evolve and affect agent performance as agentic AI coding tools mature.
+These findings establish an empirical baseline for understanding how developers configure agentic tools, suggest that AGENTS.md serves as a natural starting point, and motivate longitudinal and experimental research on how configuration strategies evolve and affect agent performance.
 Software Engineering, Generative AI, AI Agents, Configuration
@@
 Agentic AI coding tools based on large language models (LLMs) (Sergeyuk et al., 2025; Terragni et al., 2025), automate time-consuming and repetitive tasks, such as generating and editing code, tests, and documentation.
-The agentic capabilities of these tools go beyond those of purely reactive conversational assistants.
-Instead of just following user commands, these tools take initiative and act proactively to accomplish defined objectives (Anthropic, 2025).
-They can autonomously interact with development environments and use project artifacts, external data, and command-line tools to perform actions with reduced human interaction (Yang et al., 2024).
-Certain functionality can be delegated to “tools” and “(sub)agents.” Tools are capabilities with a specific bounded function (e.g., run_python(code) or search_web(query)).
-[^1] We distinguish such “tools” from full-fledged “agentic tools” (e.g., Claude Code).
-They are deterministic and invoked by the underlying model or agents.
-(Sub)agents are goal-directed action-taking components.
-They interpret a user goal, break it down into substeps, choose appropriate tools, execute actions, observe results, and potentially adjust their plan for the next iteration of this “agent loop.” Initially, agentic AI coding tools implemented one central agent loop steered by a foundational model such as Anthropic’s Opus or OpenAI’ GPT.
-Example tools include Claude Code and OpenAI Codex.
-Soon, conversational tools such as GitHub Copilot and Cursor offered similar capabilities in an “agent mode.” Over the months preceding the writing of this paper, tool vendors implemented extension and configuration mechanisms that allow developers, for example, to write their own (sub)agents that operate in parallel to the central agent loop.
-We define configuration mechanisms as means by which developers can tailor tool and agent behavior to a project or workflow (e.g., by providing contextual information in a file or by specifying dedicated subagents).
-A configuration artifact is the tangible specification of that mechanism (e.g., the actual Markdown file with the contextual information).
+Unlike purely reactive conversational assistants, these tools proactively accomplish defined objectives (Anthropic, 2025) by autonomously interacting with development environments, project artifacts, external data, and command-line tools with reduced human interaction (Yang et al., 2024).
+Certain functionality can be delegated to tools and (sub)agents.
+Tools are deterministic capabilities with a specific bounded function (e.g., run_python(code) or search_web(query)), invoked by the underlying model.
+We distinguish such fine-grained tools from full-fledged agentic tools (e.g., Claude Code).
+(Sub)agents are goal-directed components that interpret a user goal, decompose it into substeps, select and execute tools, and iteratively adjust their plan in an “agent loop.” Initially, agentic AI coding tools such as Claude Code and OpenAI Codex implemented one central agent loop steered by a foundation model, followed by conversational tools such as GitHub Copilot and Cursor offering similar capabilities in an “agent mode.” More recently, tool vendors introduced extension and configuration mechanisms that allow developers to write their own (sub)agents that operate in parallel to the central agent loop.
+We define a configuration mechanism as a means for developers to tailor tool and agent behavior to a project or workflow (e.g., context files or dedicated subagents).
+A configuration artifact is a tangible instance of a mechanism (e.g., the actual Markdown file).
 For example, files like AGENTS.md or CLAUDE.md act as “READMEs for agents” with context-specific information about build commands, coding conventions, and rules for CI/CD pipelines (Mohsenimofidi et al., 2026; OpenAI, 2025).
@@
 With this paper, we provide the following contributions: (1) We systematically documented eight configuration mechanisms: Context Files, Skills, Subagents, Commands, Rules, Settings, Hooks, and MCP servers.
-We identified these from the documentation of Claude Code, GitHub Copilot, Cursor CLI, Gemini CLI and Codex CLI.
+We identified these from the documentation of Claude Code, GitHub Copilot, Cursor CLI, Gemini CLI, and Codex CLI.
 Some mechanisms are available in all tools; others are specific to particular tools.
-(2) We analyzed the adoption of these configuration mechanisms in 2,926 GitHub repositories.
+(2) We analyzed the adoption of these configuration mechanisms in 2,923 GitHub repositories.
 Context Files (Markdown files that provide contextual project information) dominated and are often the sole configuration mechanism.
@@
 Although Skills and Subagents can be defined for a wide range of purposes, most repositories that adopt them define only one or two artifacts.
-Moreover, Skills primarily rely on static resources rather than executable scripts to extend tool and agent behaviour.
+Moreover, Skills primarily rely on static resources rather than executable scripts to extend tool and agent behavior.
 Figure 1. Data collection process.
@@
 2. Related Work
-Related work has examined how repository-level context files, prompts, and structured context mechanisms shape agent behavior, efficiency, and integration into development workflows.
-Although several studies have investigated individual configuration artifacts, most notably repository-level context files such as AGENTS.md, prior work has typically focused on single artifact types in isolation.
+Prior work has examined how repository-level context files, prompts, and structured context mechanisms shape agent behavior, efficiency, and integration into development workflows, but has typically focused on single artifact types in isolation.
 There has been no systematic examination of the broader configuration mechanisms to customize agentic behavior.
@@
 Mohsenimofidi et al.
-(Mohsenimofidi et al., 2026) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level instruction files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions.
-Complementing this work, Chatlatanagulchai et al.
-(Chatlatanagulchai et al., 2025) found that context files are actively maintained, structurally consistent, and focus predominantly on functional development instructions such as build and testing procedures.
+(Mohsenimofidi et al., 2026) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level context files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions.
+Chatlatanagulchai et al.
+(Chatlatanagulchai et al., 2025) found that context files are actively maintained through frequent, small changes and focus predominantly on development instructions such as build and testing procedures.
 Santos et al.
-(Santos et al., 2025) analyzed Claude Code projects and how configuration artifacts structure agent interactions in repositories.
+(Santos et al., 2025) analyzed 328 CLAUDE.md files and found that architectural concerns dominate, often co-occurring with development guidelines, testing, and project overview.
 Beyond structural characterization, recent work has begun to evaluate the impact of such artifacts.
@@
 (Lulla et al., 2026) conducted a controlled study comparing agent executions with and without an AGENTS.md file and reported lower runtime and token consumption while maintaining comparable task completion behavior.
-In other work, Villamizar et al.
-(Villamizar et al., 2025) argued that prompts should be treated as software engineering artifacts and outlined a research agenda to understand their evolution, reuse, and governance.
-Beyond static prompts, Zhang et al.
-(Zhang et al., 2025) introduced a framework that treats contexts as evolving instructions and addresses issues such as brevity bias and context collapse through incremental updates.
+Villamizar et al.
+(Villamizar et al., 2025) argued that prompts should be treated as software engineering artifacts and outlined a research agenda for their evolution, reuse, and governance.
+Zhang et al.
+(Zhang et al., 2025) introduced a framework that treats contexts as evolving instructions and addresses brevity bias and context collapse via incremental updates.
 Recent work further generalized this perspective by formalizing context as a programmable abstraction.
@@
 (Ye et al., 2026) proposed Meta Context Engineering, a framework in which agentic skills and context artifacts co-evolve via evolutionary search.
-However, a cross-tool and cross-agent perspective mapping configuration mechanisms and studying their adoption patterns across repositories is missing. Our work addresses this gap.
+Jiang and Nam (Jiang and Nam, 2026) explored Cursor rules as one particular configuration mechanism and developed a taxonomy of project contexts.
+However, a cross-tool perspective mapping configuration mechanisms and studying their adoption patterns across repositories is missing. Our work addresses this gap.
 3. Data Collection and Analysis
-We collected OSS projects hosted on GitHub.
-Since GitHub hosts a variety of content, we selected repositories belonging to “engineered” software projects (Munaiah et al., 2017).
-We adopted and updated a selection approach used in prior work, starting with the SEART GitHub search tool  (Mohsenimofidi et al., 2026; SEART, 2025; Dabic et al., 2021).
-We selected non-fork repositories with at least two contributors and a license that were created before 1 January 2024 with commits since 1 June 2025.
-We then excluded archived, disabled, or locked repositories We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and applied filters based on commit and watcher count based on prior research (Mohsenimofidi et al., 2026).
-This resulted in a final sample of 37,249 repositories.
+We collected OSS projects hosted on GitHub, selecting repositories belonging to “engineered” software projects (Munaiah et al., 2017) using an updated version of a selection approach from prior work, starting with the SEART GitHub search tool (Mohsenimofidi et al., 2026; SEART, 2025; Dabic et al., 2021).
+To ensure sufficient project maturity and development history, we selected non-fork repositories with at least two contributors and a license that were created before 1 January 2024 with commits since 1 June 2025.
+We then excluded archived, disabled, or locked repositories.
+We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and applied filters based on commit and watcher count based on prior research (Mohsenimofidi et al., 2026).
+This resulted in a sample of 37,249 repositories.
 Table 1.
@@
 MCP External tool or data connections via the Model Context Protocol. .mcp.json b .codex/ config.toml .cursor/ mcp.json .gemini/ settings.json
-a Copilot also supports CLAUDE.md, AGENTS.md, and GEMINI.md; b Configured via the GitHub web interface, not via files in the project repository;
+a Copilot also supports CLAUDE.md, AGENTS.md, and GEMINI.md; b Configured via the GitHub web UI, not via files in the project repository;
 c Cursor deprecated .cursorrules and now suggests using AGENTS.md instead.
-Table 2. Release dates of agentic AI coding tools and repositories ($n=2,926$; one repository can use multiple tools).
+Table 2. Release dates of agentic AI coding tools and repositories ($n=2,923$; one repository can use multiple tools).
 Agentic Tool Release (Month/Year) #Repositories
-Claude Code 02/2025 (CLI & agents since release) 1,310
+Claude Code 02/2025 (CLI & agents since release) 1,305
 GitHub Copilot 10/2021 (Release)
@@
 09/2025 (Copilot CLI)
-958
-Codex CLI 04/2025 (CLI & agents since release) 558
+957
+AGENTS.mdd 05/2025 (Adoption in Codex)
+08/2025 (Specification (agentsmd community, 2025))
+528
 Cursor CLI 03/2023 (Release)
@@
 08/2025 (Cursor CLI)
-355
+327
 Gemini CLI 02/2024 (Release)
@@
 06/2025 (Gemini CLI)
-186
-Table 3. Repository metadata comparison by agentic tool. Cells show median, IQR, and Cliff’s $\delta$ for tool-adopting repositories. Significant differences (BH-adjusted) are in bold.
+178
+Codex CLI 04/2025 (CLI & agents since release) 4
+d Repositories using AGENTS.md without any tool-specific configuration artifact.
+Table 3.
+Repository metadata by agentic tool.
+Cells show median, IQR, and Cliff’s $\delta$ for tool-adopting repositories, each compared against the complement of non-adopting repositories.
+Significant differences (BH-adjusted) are in bold.
 Agentic tool Age (years) Contrib. Commits Size (KB)
-All 6.7 42 2,106 40k
-(nn=2,926) (4.3–9.4) (19–105) (965–5,064) (9,498–132k)
-Claude 6.2∗∗∗ 40 2,234 42k
-(nn=1,310) (4.1–9.2) (19–111) (995–5,268) (10k–151k)
-$\delta$=-0.08 $\delta$=0.00 $\delta$=0.03 $\delta$=0.04
-Gemini 6.8 66∗∗∗ 3,322∗∗∗ 58k∗∗
-(nn=186) (4.7–9.5) (32–144) (1,290–9,776) (15k–193k)
-$\delta$=0.03 $\delta$=0.21 $\delta$=0.21 $\delta$=0.12
-Codex 6.9 43 1,809∗∗ 34k∗∗
-(nn=568) (4.2–9.3) (19–89) (877–4,058) (8,696–94k)
-$\delta$=0.01 $\delta$=-0.02 $\delta$=-0.10 $\delta$=-0.08
-Copilot 7.1∗∗∗ 42 2,200 44k
-(nn=958) (5.0–9.7) (19–113) (1,000–5,591) (9,858–143k)
-$\delta$=0.11 $\delta$=0.02 $\delta$=0.04 $\delta$=0.03
-Cursor 5.5∗∗∗ 54∗∗ 2,780∗∗∗ 75k∗∗∗
-(nn=355) (3.4–7.9) (23–120) (1,264–5,968) (21k–201k)
-$\delta$=-0.19 $\delta$=0.10 $\delta$=0.14 $\delta$=0.22
-Note: Mann–Whitney U test (two-sided) with Benjamini–Hochberg FDR correction (20 comparisons). Effect sizes: Cliff’s delta. $p<.05$; ∗∗ $p<.01$; ∗∗∗ $p<.001$ (adjusted).
-Figure 2. Adoption of agentic tools per programming language; % relative to #repositories using a certain tool.
-Figure 3. Usage of configuration mechanisms across agentic tools. % relative to #repositories using a certain mechanism.
+All 6.7 42 2,108 39k
+(nn=2,923) (4.3–9.4) (19–105) (965–5,065) (9,501–132k)
+Claude 6.2∗∗∗ 40 2,217 42k
+(nn=1,305) (4.1–9.2) (19–111) (989–5,270) (9,979–152k)
+$\delta$=-0.08 $\delta$=-0.00 $\delta$=0.03 $\delta$=0.04
+Copilot 7.1∗∗∗ 42 2,231 45k
+(nn=957) (5.0–9.7) (19–113) (1,001–5,607) (9,958–143k)
+$\delta$=0.11 $\delta$=0.03 $\delta$=0.04 $\delta$=0.04
+AGENTS.md 7.1 44 1,809∗∗∗ 31k∗∗∗
+(nn=528) (4.4–9.4) (19–89) (869–3,972) (8,511–91k)
+$\delta$=0.03 $\delta$=-0.02 $\delta$=-0.10 $\delta$=-0.10
+Cursor 5.5∗∗∗ 52 2,780∗∗∗ 75k∗∗∗
+(nn=327) (3.4–7.9) (23–118) (1,264–6,029) (21k–198k)
+$\delta$=-0.20 $\delta$=0.09 $\delta$=0.14 $\delta$=0.22
+Gemini 6.8 63∗∗∗ 3,302∗∗∗ 58k
+(nn=178) (4.7–9.5) (31–142) (1,282–9,711) (15k–189k)
+$\delta$=0.02 $\delta$=0.20 $\delta$=0.20 $\delta$=0.12
+Codex 7.6 46 6,014 505k
+(nn=4) (4.6–11) (32–62) (4,997–6,514) (153k–939k)
+$\delta$=0.12 $\delta$=-0.07 $\delta$=0.48 $\delta$=0.66
+Note: Mann–Whitney U test (two-sided) comparing each tool’s adopters against non-adopters (complement), with Benjamini–Hochberg FDR correction (6 categories $\times$ 4 metrics = 24 comparisons).
+Effect sizes: Cliff’s delta.
+$p<.05$; ∗∗ $p<.01$; ∗∗∗ $p<.001$ (adjusted).
+Interpret Codex results (nn=4) with caution due to sample size.
+Figure 2.
+Adoption of agentic tools per programming language; AGENTS.md denotes repositories using only that file with no tool-specific configuration.
+Percentages relative to repository count per tool.
+Column totals can exceed the overall repository count (multiple tools per repository).
 Diagram showing adoption of AI coding tools per programming language.
-Figure 1 outlines our data collection pipeline for these repositories.
-We cloned them and searched their default branch for a README file.
-We excluded 11 repositories that became unavailable in the meantime, 60 repositories without a README file, and then used the lingua-language-detector Python library to detect the language of files we found.
-This resulted in 994 repositories with non-English README files being excluded.
-For the remaining 36,184 repositories, we developed a classification pipeline that uses the GPT-5.2 model to determine, based on the content of a README file, whether the software project fulfills our definition of an “engineered” project.
-We based our definition on Munaiah et al.’s paper (Munaiah et al., 2017).
-An engineered project is a project that shows clear evidence of software engineering practices (e.g., it includes a clearly stated purpose of a problem that the software solves; it provides installation and deployment instructions, developer and user documentation, testing, CI/CD, or quality-assurance-related information).
-There should also be evidence for maintenance.
-This definition was part of our prompt, together with the README of each repository.
-In our supplementary material, we share the final prompt and previous iterations of the prompt, which we tested with subsets of the data, as well as the exact configuration for labeling repositories as “engineered” or not (we kept OpenAI’s default values for all parameters) (Authors, 2026).
-The classification pipeline labeled 32,564 repositories as engineered software projects, was ‘unsure’ about 2,204 repositories, and excluded 1,416 repositories.
-For each label category, we spot-checked randomly selected samples, both during the iterative development of the prompts and for the final sample.
+Figure 3.
+Usage of configuration mechanisms across agentic tools.
+Repositories using only AGENTS.md without tool-specific configuration are excluded.
+Percentages relative to repository count per tool.
+Column totals can exceed the overall repository count (multiple tools per repository).
+Diagram showing usage of configuration mechanisms across agentic tools.
+Figure 1 outlines the data collection pipeline.
+We cloned the repositories and searched their default branch for a README file, excluding 11 that became unavailable and 60 without a README.
+We then used the lingua-language-detector Python library to detect each file’s language, excluding 994 repositories with non-English README files.
+For the remaining 36,184 repositories, we used a GPT-5.2-based classification pipeline to determine, based on README content, whether each project qualifies as “engineered” per Munaiah et al.
+(Munaiah et al., 2017), that is, showing clear evidence of software engineering practices (e.g., a stated purpose, installation and deployment instructions, documentation, testing, CI/CD) and maintenance.
+The supplementary material includes the prompt (iteratively developed and tested on subsets) and model configuration (OpenAI defaults) (Authors, 2026).
+The pipeline labeled 32,564 repositories as engineered and excluded the remaining 3,620 (1,416 classified as non-engineered, 2,204 “unsure”).
+We spot-checked randomly selected samples from each category during prompt development and for the final sample.
 We then cloned the remaining 32,564 repositories and applied heuristics based on file names and file paths to detect the usage of AI coding tools and configuration mechanisms (see Table 1).
 These heuristics are based on the answers to RQ1 and are briefly discussed in Section 4.
-This resulted in 2,926 repositories that use one or more AI coding tools.
+This resulted in 2,923 repositories that use one or more AI coding tools (three additional repositories contained only empty configuration artifacts and were excluded).
 Our data collection and analysis scripts and the analyzed data are available online (Authors, 2026).
 4. Configuration Mechanisms (RQ1)
-To determine the set of configuration mechanisms for agentic AI coding tools, we followed two steps.
-We first selected five agentic AI coding tools based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026).
-The tools represent the four most popular AI tools among the surveyed developers.
-We added Codex instead of ChatGPT as Codex is an agentic tool by the same vendor using the same models.
-We added Cursor because it has recently been studied in software engineering  (He et al., 2026; Jiang and Nam, 2026).
-Table 2 lists the five tools.
-Our study focuses on agentic tools with a command-line interface (CLI) that first appeared in 2025.
-For some tools, prior non-agentic versions exist.
-In the second step, one author systematically reviewed the online documentation of each tool, documenting configuration mechanisms together with repository-level files and directories that might indicate their usage.
-This was then cross-checked by two other authors and discussed among all.
-As part of RQ1 and to “operationalize” configuration mechanisms and their occurrence in repositories, we then used that information to develop heuristics to identify whether tools are used in a repository (see Table 1).
-Note that for GitHub Copilot, Cursor, and Gemini, files and directories indicating their use apply both to their conversational and agentic interfaces.
-For example, a .claude folder or a CLAUDE.md file indicate usage of Claude Code.
-In addition, we developed strategies to determine whether configuration mechanisms were used.
-For example, if a directory .claude/agents/ with Markdown files is present, this indicates usage of Subagents.
+We selected five agentic AI coding tools (Table 2) based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026): the four most popular AI tools among the surveyed developers, substituting Codex for ChatGPT as the agentic tool by the same vendor, plus Cursor, which has recently been studied in software engineering (He et al., 2026; Jiang and Nam, 2026).
+Our study focuses on the CLI-based agentic interfaces that first appeared in 2025; for some tools, prior non-agentic versions exist.
+One author systematically reviewed each tool’s online documentation, documenting configuration mechanisms together with repository-level files and directories that indicate their usage.
+This was then cross-checked by two other authors and discussed among all three.
+We used these findings to develop heuristics for detecting tool usage and configuration mechanisms in repositories (Table 1).
+For example, a .claude directory or a CLAUDE.md file indicates Claude Code usage; a .claude/agents/ directory with Markdown files indicates Subagents.
+AGENTS.md is a special case: as a tool-agnostic standard, it is supported by multiple tools, so repositories using only AGENTS.md without tool-specific artifacts are tracked separately (see Table 2, footnote d).
+Note that for Copilot, Cursor, and Gemini, certain detected files apply to both conversational and agentic interfaces.
 All documents and the Python scripts that implement our matching strategy are part of the supplementary material (Authors, 2026).
 {results}
-Key insights for RQ1:
-  • Qualitative analysis of the documentation of agentic tools revealed eight configuration mechanisms.
-  • There is convergence for some mechanisms across tools (e.g., all five support Context Files and Skills), yet no single tool implements all eight mechanisms, suggesting the ecosystem is still maturing, standards are developing.
+RQ1 (Summary):
+  • We identified eight configuration mechanisms spanning a spectrum from static context (e.g., Context Files) to executable and external integrations (e.g., Skills, MCP).
+Two mechanisms (Context Files and Skills) are supported by all five tools.
+  • Despite this convergence, no single tool implements all eight mechanisms.
 5. Adoption of Configuration Mechanisms (RQ2)
@@
 5.1. Characterization of Repositories in Dataset
-Before analyzing adoption, we characterize the 2,926 repositories in our dataset to contextualize the configuration mechanisms of agentic AI coding tools.
-Most repositories adopted a single tool.
+Before analyzing adoption, we characterize the 2,923 repositories in our dataset to contextualize the subsequent analysis.
+Of the 2,923 repositories, 2,018 (69.0%) adopted a single tool, while 299 (10.2%) configured exactly two tools and 50 (1.7%) configured three or more.
+An additional 528 repositories (18.1%) use AGENTS.md as a tool-agnostic standard without any tool-specific configuration artifact; these are shown separately in the analysis as “AGENTS.md” (see Table 2, footnote d).
 Among repositories with multiple tools, Claude appeared most frequently with others.
-The most common combination was Claude and Copilot ($n=128$).
-Overlaps of three or more tools were rare.
-In our initial sample, the top five main programming languages per repository ($n=36,194$) were Python (8,133; 22.5%); TypeScript (4,999; 13.8%); Java (3,813; 10.5%); Go (3,786; 10.5%); and JavaScript (3,709; 10.3%).
-Interestingly, this order is slightly different for repositories that use agentic coding tools ($n=2,926$, see Figure 3) where TypeScript and Go were more prominent and C# replaces JavaScript: TypeScript (922, 31.5%); Python (548, 18.7%); Go (533, 18.2%); C# (256, 8.8%); Java (246, 8.4%).
-TypeScript is the most common primary language for repositories using Claude, Codex, Copilot, and Cursor.
+The most common combination was Claude and Copilot ($n=168$), followed by Claude and Cursor ($n=145$), Claude and Gemini ($n=54$), and Copilot and Cursor ($n=49$).
+Notably, 44.3% of Cursor repositories also configure Claude.
+The dominance of single-tool repositories limits the potential for multi-tool confounding in tool-specific analyses.
+In our initial sample, the top five main programming languages per repository ($n=36,184$) were Python (8,133; 22.5%); TypeScript (4,999; 13.8%); Java (3,813; 10.5%); Go (3,786; 10.5%); and JavaScript (3,709; 10.3%).
+Interestingly, this order is slightly different for repositories that use agentic coding tools ($n=2,923$, see Figure 2) where TypeScript and Go were more prominent and C# replaces JavaScript in the top five: TypeScript (745, 25.5%); Python (485, 16.6%); Go (457, 15.6%); C# (228, 7.8%); Java (227, 7.8%).
+TypeScript is the most common primary language for repositories using Claude, Copilot, and Cursor.
 Usage in Java and C# repositories is lower for all tools but Copilot.
-As shown in Table 3, repositories using Cursor are younger compared to those using other tools.
-Repositories associated with Gemini exhibit larger contributor counts and commit volumes, compared to those associated with other tools.
-Cursor repositories are the largest in terms of kilobytes of source code.
+As shown in Table 3, which compares each tool’s adopting repositories against the complement of non-adopting repositories, repositories using Cursor are younger compared to those using other tools.
+Excluding Codex (nn=4), repositories associated with Gemini exhibit larger contributor counts and commit volumes, compared to those associated with other tools.
+Cursor repositories are the largest in terms of kilobytes of source code (excluding Codex).
+AGENTS.md repositories (those using only AGENTS.md with no tool-specific configuration) tend to have fewer commits and smaller size than the overall population, suggesting that these repositories adopted AGENTS.md as a lightweight, tool-agnostic starting point.
 5.2. Distribution of Configuration Mechanisms
-Figure 3 shows the distribution of configuration mechanisms.
-Across repositories, Context Files (with artifacts such as CLAUDE.md or AGENTS.md files) were the most frequently adopted configuration mechanism.
-In contrast, other configuration mechanisms, including Rules, Settings, Commands, Skills, and Subagents, were substantially less prevalent.
-Each of these mechanisms appeared in fewer than 20% of the repositories.
-Rules are primarily concentrated in repositories using Cursor, since Cursor was one of the first to introduce Rules (Jiang and Nam, 2026).
-Most repositories adopt a limited subset of available configuration mechanisms.
-In many cases, repositories include only a single Context File artifact.
-The adoption of multiple non-context file configuration mechanisms within the same repository remains comparatively rare.
-These results suggest that configuration usage is currently characterized by one dominant baseline configuration (i.e., Context Files).
-Figure 4. Cumulative adoption of selected configuration artifacts for agentic tools. Copilot Instructions include the two different artifact types for Context Files, see Table 1.
+Figure 3 shows the distribution of configuration mechanisms per tool.
+Context Files (e.g., CLAUDE.md, AGENTS.md) were the most frequently adopted mechanism, used by 61.5–100% of repositories across all tools.
+Beyond Context Files, distinct tool-specific profiles emerge: 72.8% of Cursor repositories adopt Rules, since Cursor was one of the first to introduce this mechanism (Jiang and Nam, 2026), and 61.2% of Gemini repositories use Settings.
+Other mechanisms remain below 20% of repositories for most tools.
+Most repositories included only a single Context File artifact, and the adoption of multiple non-context file mechanisms within the same repository remained rare.
+These results suggest that configuration usage is characterized by one dominant baseline configuration (i.e., Context Files).
+Figure 4. Cumulative adoption of selected configuration artifacts for agentic tools. Copilot Instructions comprise two Context File artifact types (see Table 1).
+Diagram showing the cumulative adoption of configuration artifacts for agentic tools.
 Figure 5. Configuration mechanism count per repository.
-Diagram showing the comulative adoption of configuraiton artifacts for agentic tools.
+Diagram showing the configuration mechanism count per repository.
 5.2.1. Co-occurrence of Configuration Mechanisms
-Several configuration mechanisms are frequently adopted together.
-Some of these correlations can be explained by shared configuration artifacts.
-For example, Settings and Hooks have a strong positive correlation ($\rho=0.52$), because hooks are usually defined in Settings files.
-However, Settings and Hooks also both co-occur with Skills ($\rho=0.23$ and $\rho=0.24$, respectively), which are defined and configured differently.
-Subagents also show positive correlations with multiple mechanisms, including Settings, Skills, and MCP.
+We quantified pairwise associations between mechanism adoption using Cramér’s $V$ from chi-squared tests (28 comparisons, Benjamini–Hochberg corrected).
+Several configuration mechanisms are frequently adopted together, and some of these associations can be explained by shared configuration artifacts.
+For example, Settings and Hooks show a positive association ($V=0.36$), because hooks are usually defined in Settings files.
+Settings and Hooks also both co-occur with Skills ($V=0.15$ and $V=0.24$, respectively), which are defined and configured differently.
+Subagents also show positive associations with multiple mechanisms, including Settings, Skills, and MCP.
 Other configuration mechanisms appear less frequently in combination.
-Context Files are negatively correlated with Rules ($\rho=-0.36$) and Commands ($\rho=-0.14$), which can again attributed to how they are defined and which tools first introduced the mechanisms.
-Overall, configuration mechanism adoption across repositories is characterized by recurring combinations of mechanisms rather than uniform usage.
-Patters of co-occurrence might be impacted by the tools and configuration mechanisms they support (e.g., Claude, the dominating tool, does not support Rules).
+Context Files are negatively associated with Rules ($V=0.41$) and Commands ($V=0.07$), which can be attributed to how they are defined and which tools first introduced the mechanisms.
+Overall, configuration mechanism adoption is characterized by recurring combinations rather than uniform usage, shaped by the tools and mechanisms they support (e.g., Claude, the dominating tool, does not support Rules).
+The full co-occurrence matrix is available in the supplementary material.
 5.2.2. Adoption of Configuration Mechanisms over Time
-In Figure 5, we show the adoption of configuration mechanisms over time.
-The figure shows individual artifacts that together form the mechanism Context Files, plus Skills and Subagents which have a uniform format.
-Context Files (CLAUDE.md, AGENTS.md, copilot-instructions.md) clearly dominate and increase continuously, while Skills and Subagents experienced a comparatively slow growth.
-As the figure shows, .cursorrules and copilot-instructions.md started being introduced in 2024.
-Although Cursor and also Copilot were originally released in 2023 and 2021, respectively, agentic capabilities were only introduced in 2025 (see Table 2).
-The adoption of copilot-insturctions.md has increased since then.
+Figure 4 shows the adoption over time of individual Context File artifacts, plus Skills and Subagents.
+Context Files (e.g., CLAUDE.md, AGENTS.md) clearly dominated and increased continuously, while Skills and Subagents experienced comparatively slow growth.
+.cursorrules and copilot-instructions.md started being introduced in 2024, although Cursor and Copilot were originally released in 2023 and 2021, respectively; their agentic capabilities were only introduced in 2025 (see Table 2).
+The adoption of copilot-instructions.md has increased since then.
 {results}
-Key insights for RQ2:
-  • Context Files are the dominant and often sole configuration mechanism. Advanced configuration mechanisms (e.g., Hooks, MCP) are adopted less frequently.
-  • Adoption varies: Claude has a broad configuration footprint, while Cursor emphasizes rule-based mechanisms.
-  • Although certain correlations between configuration mechanisms exist, the landscape still appears in flux.
+RQ2 (Summary):
+  • Context Files are the dominant and often sole configuration mechanism; other mechanisms are adopted less frequently.
+Over time, the gap has widened, with Skills and Subagents experiencing comparatively slow growth.
+  • Adoption varies: Claude has a broad configuration footprint, while Cursor emphasizes rule-based mechanisms. Co-occurrence patterns are shaped by tool-specific mechanism support.
+  • Most repositories (69.0%) adopt a single tool, and multi-tool overlap is limited. An additional 18.1% use AGENTS.md without any tool-specific configuration.
 6. Details of Configuration Mechanisms (RQ3)
-In this section, we analyze the use of three configuration mechanisms in more detail.
-We selected Context Files and Skills because they are supported by all tools, as discussed in Section 5, and Subagents because they follow a similar format as Skills.
+We analyzed three mechanisms in detail: Context Files and Skills because they are supported by all tools, and Subagents because they follow a similar format as Skills.
 6.1. Configuration Mechanism: Context Files
-Context Files are Markdown files that provide a central machine-readable source of contextual information about the repository in which they are used.
-AGENTS.md was introduced by OpenAI and now serves as an open tool-agnostic convention that is increasingly supported by various tools (Mohsenimofidi et al., 2026).
-We analyzed the presence and co-occurrence of Context Files across repositories.
-We identified 4,860 Context Files across 2,634 of the 2,926 repositories in our sample.
-The histogram in Figure 5 shows the distribution of Context Files.
-Most repositories that do use Context Files include one or two of such files.
-CLAUDE.md emerges as the dominant file type with 1,661 (34.2%), followed closely by AGENTS.md and copilot-instructions.md with 1,572 (32.3%) and 1,344 (27.7%) files, respectively.
+Context Files are Markdown files that provide central, machine-readable context about a project.
+AGENTS.md, introduced by OpenAI, now serves as an open tool-agnostic convention increasingly supported by various tools (Mohsenimofidi et al., 2026).
+We identified 4,854 Context Files across 2,631 of the 2,923 repositories in our sample.
+Most repositories include one or two such files (Figure 5).
+CLAUDE.md emerges as the dominant file type with 1,658 (34.2%), followed closely by AGENTS.md and copilot-instructions.md with 1,571 (32.4%) and 1,393 (28.7%) files, respectively.
 GEMINI.md (159 files, 3.3%) and .cursorrules (73 files, 1.5%) are rare.
@@
 Cursor rules are different and are stored in .cursor/rules/.
-We also investigated the repository-level adoption of the Context Files, where CLAUDE.md demonstrates the highest adoption rate with 45.40% of repositories (1,195 repos), AGENTS.md follows closely with 40.6% adoption (1,069 repos), and copilot-instructions.md appears in 35.1% of repositories (925 repos).
-The predominance of CLAUDE.md, AGENTS.md and copilot-instructions.md, both in terms of their frequency and the number of repositories adopting them, indicates a clear concentration of projects around these files.
-We then investigated programming language-specific patterns in Context Files usage across the repositories.
-We found Context Files in 651 repositories with TypeScript as the main language, 435 with Python, 414 with Go, 210 with Java, 210 with C#, 201 with JavaScript, 134 with Ruby, 181 with Rust, 156 with C++, 123 with PHP, and 53 with C.
-CLAUDE.md is the most common Context File across languages, except for Java, C#, and C++, where copilot-instructions.md emerges as the dominant type.
+We also investigated the repository-level adoption of the Context Files: among the 2,631 repositories with Context Files, CLAUDE.md demonstrates the highest adoption rate with 45.4% of repositories (1,195 repos), AGENTS.md follows closely with 40.6% adoption (1,068 repos), and copilot-instructions.md appears in 35.4% of repositories (930 repos).
+Examining language-specific patterns, Context Files appeared most frequently in TypeScript (651 repos), Python (435), Go (414), C# (210), Java (209), JavaScript (200), Rust (180), C++ (156), PHP (123), and C (53) repositories.
+CLAUDE.md was the most common Context File across languages, except for Java, C#, and C++, where copilot-instructions.md dominated.
 Figure 6. Creation order of Context Files per repository. Curly braces indicate that files were added on the same day.
+Diagram showing the creation order of Context Files per repository.
 Figure 7. Reference-only Context Files. Examples of other files include README.md and CONTRIBUTING.md.
-Diagram showing the creation order of Context Files per repository.
-Figure 7 shows the creation order of the different Context Files.
-Most repositories only used one type of Context Files.
-These repositories are not captured in this figure.
-AGENTS.md and CLAUDE.md are typically created first, and if CLAUDE.md existed first, then AGENTS.md was commonly created afterward.
-A reason for this could be that Claude Code is the most popular agentic AI coding tool, but does not yet support the emerging standard AGENTS.md (anthropics/claude-code on GitHub, 2026).
-Interestingly, repositories that started with copilot-instructions.md then often created CLAUDE.md or AGENTS.md files, even though GitHub Copilot supports all major Context Files types.
-This, in addition to the overall dominance of CLAUDE.md and AGENTS.md, illustrates that these two files currently serve as the de facto standard.
-Finally, we observed that Context Files commonly refer to each other.
-This is visualized in Figure 7.
-“Referring to” means that one file, the source, contains a reference to another file, the target, rather than providing their own content.
-We found three types of references between Context Files:
+Diagram showing the reference patterns between Context Files.
+Figure 6 shows the creation order of Context Files in repositories with multiple types (single-type repositories are omitted).
+CLAUDE.md was typically created first, with AGENTS.md commonly added afterward—likely because Claude Code is the most popular agentic tool but does not yet support the emerging standard AGENTS.md (anthropics/claude-code on GitHub, 2026).
+Repositories that started with copilot-instructions.md often later added CLAUDE.md or AGENTS.md, despite Copilot supporting all major Context File types, reinforcing their de facto standard status.
+Finally, Context Files commonly reference each other (Figure 7). We identified three types of references, where a source file points to a target file rather than providing its own content:
   1. (1)
@@
     Brief summary plus reference: Context Files include a header, 1–2 sentences of context, and a reference to another file, e.g., # AI Guidelines […] **Read @AGENTS.md for comprehensive guidelines.**
-We found 518 file pairs that reference each other.
-The file type with most references was CLAUDE.md, with 357 “outgoing” references, followed by AGENTS.md and GEMINI.md.
-In contrast, the most frequently referenced file was AGENTS.md with 368 “incoming” references, substantially more than CLAUDE.md and copilot-instructions.md.
-Furthermore, the strongest pattern was the CLAUDE.md to AGENTS.md pair that occurred 311 times.
-In contrast, when AGENTS.md files acted as a pointer, they most frequently referenced CLAUDE.md files.
-This again reflects the symbiotic relationship between these two types of Context Files that we also found in the creation order discussed earlier.
+We found 518 reference pairs.
+CLAUDE.md had the most outgoing references (357), predominantly pointing to AGENTS.md (311 times).
+Conversely, AGENTS.md received the most incoming references (368) and, when acting as a pointer, most frequently referenced CLAUDE.md—reflecting the symbiotic relationship also observed in the creation order.
 6.2. Configuration Mechanism: Skills
-Skills were introduced by Anthropic and now serve as an open standard for extending the capabilities of agentic AI tools with specialized contextual knowledge and workflows.
+Skills were introduced by Anthropic and now serve as an open standard for extending agentic AI tools with specialized contextual knowledge and workflows.
 Skills bundle prompts, tools, and documentation that an agent can invoke on demand.
-From an implementation perspective, Skills are a directory that contains a SKILL.md file.
-This file must include a YAML frontmatter with at least a name and description, followed by instructions in the body for how a task should be performed.
+Each Skill is a directory containing a SKILL.md file with YAML frontmatter (name, description) and instructions in the body.
 We found 601 Skills in 158 repositories.
@@
 However, as Figure 5 shows, repositories that use Skills typically use fewer than three Skills.
-As with Context Files, we also explored whether there are programming language-specific patterns.
-Figure 3 shows that TypeScript and Python are the dominating programming languages in all repositories that use agentic tools.
-This distribution is also reflected across projects that use Skills.
-The specification of Skills recommends that SKILL.md files should contain fewer than 500 lines.
-It further recommends moving any detailed material to separate files (agentskills.io, 2026).
-We found that of the 601 Skills we identified, only 25 (4%) were longer than 500 lines.
-That is, most developers follow the recommendation.
-To provide detailed information and instructions, Skills can utilize different resources, i.e., additional optional files or directories that are loaded by the agents of a tool when required.
-There are three types of resource:
+As with Context Files, TypeScript and Python dominated the language distribution of repositories that use Skills, mirroring their dominance across all repositories that use agentic tools (Figure 2).
+The specification recommends that SKILL.md files contain fewer than 500 lines, with detailed material moved to separate files (agentskills.io, 2026).
+Only 29 of the 601 Skills (5%) exceeded this limit.
+To provide detailed information and instructions, Skills can include additional resource directories loaded by agents when required:
   1. (1)
@@
     assets/ contains static resources, such as templates (document templates, configuration templates), images (diagrams, examples) and data files (lookup tables, schemas).
-To obtain the resources used by the different Skills, we scanned all Skills directories in the studied repositories for the occurrence of the relevant resource folders.
-Of the 601 skills analyzed, the vast majority (501, 83.3%) included no additional resources.
-Among those that did, the most common pattern was a references/ directory (48 skills, 8.0%), followed by a scripts/ directory (34, 5.7%).
-Twelve Skills (2.0%) combined both scripts/ and references/, while an assets/ directory appeared in only four Skills (0.7%).
+We scanned all Skills directories for the occurrence of resource folders.
+Of the 601 skills analyzed, the vast majority (514, 85.5%) included no additional resources.
+Among those that did, the most common patterns were a references/ directory and a scripts/ directory (35 skills each, 5.8%).
+Eleven Skills (1.8%) combined both scripts/ and references/, while an assets/ directory appeared in only four Skills (0.7%).
 The remaining combinations, i.e., references/ with assets/, and all three directories together, each occurred just once (0.2%).
@@
 6.3. Configuration Mechanism: Subagents
-Subagents are specialized AI agents that agentic AI coding tools or other agents can delegate tasks to.
-Technically, Subagents are a Markdown file with the same YAML frontmatter structure as Skills.
-However, Subagents differ in that they operate within their own context window and return the results to the parent agent whereas Skills execute instructions within the calling agent’s context (Anthropic, 2026; Cursor, 2026).
-We found 452 Subagents across 131 repositories.
-On average, these repositories use 3.45 Subagents ($min=1.0$; $max=18.0$; $median=2$).
-However, Figure 5 indicates that these repositories commonly use fewer than three Subagents.
-One feature of Claude Code’s Subagents is that they can have their own “memory”, i.e., a persistent directory that survives across interactions with the agent (Anthropic, 2026).
-This directory can be used to build up knowledge over time, e.g., to store debugging insights.
-However, we could not find any repositories storing such memory files.
+Subagents are specialized AI agents to which agentic AI coding tools or other agents can delegate tasks.
+They share the same YAML frontmatter structure as Skills but operate within their own context window and return results to the parent agent, whereas Skills execute within the calling agent’s context (Anthropic, 2026; Cursor, 2026).
+We found 450 Subagents across 131 repositories.
+On average, these repositories use 3.44 Subagents ($min=1.0$; $max=17.0$; $median=2$), though most use fewer than three (Figure 5).
+Claude Code’s Subagents can have their own “memory,” i.e., a persistent directory that survives across interactions and can build up knowledge over time, e.g., to store debugging insights (Anthropic, 2026).
+However, we found no repositories storing such memory files.
 {results}
-Key insights for RQ3:
+RQ3 (Summary):
   • Context Files: CLAUDE.md, AGENTS.md and copilot-instructions.md are the most common artifacts.
@@
   • Skills: Although Skills can be fine-grained and offer various customization features, most repositories that adopt them only define one or two and predominantly utilize static documentation rather than dynamic scripts.
-  • Subagents: The usage patterns for Subagents are similar to those of Skills. Future work should study differences in content and use cases in more detail.
+  • Subagents: Usage patterns mirror those of Skills, with most repositories defining only one or two. No repositories use the persistent memory feature available for Claude Code’s Subagents.
 6.4. Threats to Validity
@@
 We partially mitigate this by restricting our sample to actively maintained repositories, but artifact presence remains a proxy for adoption.
+As a further mitigation, we scanned the complete commit history of every repository with at least one configuration file for AI-authored commits, using author/committer identity fields, git trailers (e.g., Co-authored-by), and commit message patterns.
+Among the 2,895 repositories with at least one configuration artifact, 2,079 (71.8%) also contain AI-authored commits, providing evidence that detected artifacts are associated with active tool usage rather than being dormant files.
+The detection scripts and full results are included in the supplementary material.
 Additionally, for GitHub Copilot, Cursor, and Gemini, detected files apply to both conversational and agentic modes, so we cannot isolate agentic usage.
 This limitation does not affect agentic-only mechanisms such as Skills and Subagents, nor the broader AGENTS.md trend.
+Although multi-tool overlap could conflate tool-specific patterns, 85.3% of tool-adopting repositories configure a single tool, limiting this concern.
+To verify that detected artifacts contain meaningful content, we programmatically scanned all configuration files across the eight artifact types and excluded empty files (reducing the initial 2,926 detected repositories to 2,923).
+One of the authors then used Claude Code (Opus 4.6 with high effort) to check all 451 files with ten or fewer lines.
+Of these, 396 contained substantive configuration content (e.g., coding conventions, agent definitions, tool settings), 51 were reference-only files analyzed separately in Section 6, and four were borderline cases (e.g., a bare heading or filename).
+The supplementary material includes the inspection script, the checked files, and their classification.
 Internal validity:
@@
 7. Discussion
-Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,926 repositories.
+Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,923 repositories.
 Below, we discuss the implications of our study.
@@
 Our results reveal a clear trajectory toward AGENTS.md as a unifying, tool-agnostic configuration artifact.
-The creation order and reference patterns point in the same direction: For example, when both files are present, CLAUDE.md typically appears first and AGENTS.md is added later (Figure 7).
-CLAUDE.md most frequently points to AGENTS.md (311 cases), and AGENTS.md receives 368 incoming references overall, far more than any other file type.
-Repositories that began with copilot-instructions.md often later introduced CLAUDE.md or AGENTS.md, despite GitHub Copilot already supporting all major Context File types.
+Creation order analysis shows that CLAUDE.md typically appears first with AGENTS.md added later (Figure 6).
+Reference patterns confirm this.
+CLAUDE.md most frequently points to AGENTS.md (311 cases), which receives 368 incoming references overall, far more than any other file type.
 Together, these patterns suggest bottom-up convergence around AGENTS.md as a cross-tool baseline, driven by developer practice and compatibility needs rather than vendor mandate.
+This is further reinforced by the 528 repositories that use AGENTS.md without any tool-specific configuration, suggesting that developers adopt it as a tool-agnostic standard even before committing to a specific tool.
 At the same time, the layering of multiple context files within a single repository raises the risk of redundant or conflicting instructions across artifacts.
@@
 Most repositories include only one or two Context Files (Figure 5), and the median number of Skills and Subagents per repository is two in both cases.
-Moreover, 83.3% of Skills do not include additional resources, and when resources are present, static documentation (references/) is more common than executable scripts (scripts/).
+Moreover, 85.5% of Skills do not include additional resources, and when resources are present, static documentation (references/) is equally common as executable scripts (scripts/).
 In practice, Skills, therefore, function primarily as structured text rather than executable workflow bundles.
@@
 We also found no evidence of repositories using the persistent memory feature of Claude Code’s Subagents.
-This gap between available depth and actual usage likely reflects both the novelty of these mechanisms and the effort required to configure them.
-Claude Code was released only in February 2025, with Skills and Subagents introduced even more recently.
-Beyond recency, developers may gravitate toward the lowest-friction mechanism (i.e., Context Files) without exploring more advanced options.
-Defining executable Skills with scripts and structured resources requires additional design and maintenance effort compared to authoring Markdown instructions, which may deter adoption in the absence of clear guidance on best practices.
-At present, there is little empirical evidence on which configuration strategies are most effective or under which conditions they yield measurable improvements.
-Future work should therefore assess whether deeper configuration leads to performance gains, extending early evidence on the impact of Context Files (Lulla et al., 2026).
+This gap likely reflects both the novelty of these mechanisms and the effort required to configure them.
+Developers may gravitate toward the lowest-friction mechanism (i.e., Context Files), since defining executable Skills with scripts and structured resources requires more design and maintenance effort than authoring generic Markdown-based files.
+Future work should assess whether deeper configuration leads to measurable performance gains, extending early evidence on the impact of Context Files (Lulla et al., 2026).
 Distinct tool ecosystems:
-Configuration practices vary systematically across tools.
-Claude Code repositories exhibit the broadest configuration footprint, whereas Cursor projects emphasize Rules and Commands, and Copilot and Codex repositories rarely extend beyond Context Files.
+Configuration practices vary systematically across tools: Claude Code repositories exhibit the broadest configuration footprint, Cursor projects emphasize Rules and Commands, and Copilot repositories rarely extend beyond Context Files.
 These differences likely reflect both the configuration options each tool exposes (Table 1) and the norms emerging within their respective user communities.
-Tool ecosystems also differ in repository characteristics: Cursor repositories tend to be younger and larger, while Gemini repositories show comparatively higher activity levels (Table 3).
-An open question is whether these tool-specific configuration cultures will converge as feature sets overlap, or whether distinct patterns of use will persist.
+Repository characteristics also differ: Cursor repositories tend to be younger and larger, while Gemini repositories show higher activity levels (Table 3).
+The 528 repositories using AGENTS.md without any tool-specific configuration further reinforce its role as a tool-agnostic standard.
+It remains an open question whether these tool-specific configuration cultures will converge as tool capabilities increasingly overlap, or whether distinct usage patterns will persist.
 Practical implications and future directions:
-Our systematically compiled overview of eight configuration mechanisms (Table 1) provides a reference for developers seeking to understand the options available to them.
-Based on our findings, we offer the following concrete observations for practitioners.
+Table 1 provides a reference for developers seeking to understand available configuration options.
+Based on our findings, we offer the following observations for practitioners.
 First, Context Files—particularly AGENTS.md—represent the lowest-friction entry point for configuring agentic tools and are already widely adopted.
-Second, developers who rely on multiple tools should consider maintaining an AGENTS.md file as a shared configuration baseline, given its cross-tool support and the reference patterns we observed.
-When multiple context files coexist, teams may benefit from clearly structuring them hierarchically (e.g., using tool-specific files as adapters that reference a shared core file) to reflect the layering patterns we identified.
-Third, for recurring, well-defined workflows, Skills offer the potential for richer configuration through scripts and structured resources, but this potential is currently underutilized.
-In practice, configuration is presently used more as documentation than as automation, suggesting that teams should adopt executable Skills deliberately and only when the expected benefits justify the additional setup effort.
-Finally, given the distinct configuration profiles across tools, teams should be mindful that adopting tool-specific mechanisms (e.g., rule-based configurations) may shape their workflows in tool-dependent ways.
+Second, developers who rely on multiple tools should maintain an AGENTS.md file as a shared configuration baseline, given its cross-tool support and the reference patterns we observed.
+When multiple context files coexist, teams may benefit from structuring them hierarchically (e.g., using tool-specific files as adapters that reference a shared core file).
+Third, for recurring, well-defined workflows, Skills offer potential for richer configuration through scripts and structured resources, but in practice Skills are presently used more as static documentation than as executable workflows.
+Finally, adopting tool-specific mechanisms (e.g., rule-based configurations) may shape workflows in tool-dependent ways.
 For researchers, our findings point to several directions.
-Longitudinal studies are needed to track how configuration practices evolve as tools mature and as developers gain experience.
-Controlled studies should assess whether advanced mechanisms such as Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone, extending prior work on the operational impact of AGENTS.md (Lulla et al., 2026).
-Given that projects increasingly use multiple tools (Section 5.1), research should also investigate interoperability challenges: how configuration artifacts interact across tools, and how potential conflicts between overlapping or contradictory instructions in different files can be detected and resolved.
+Longitudinal studies are needed to track how configuration practices evolve as tools mature.
+Controlled studies should assess whether advanced mechanisms such as Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone (Lulla et al., 2026).
+Given that certain projects use multiple tools (Section 5.1), research should also investigate interoperability challenges, including how conflicts between overlapping or contradictory instructions across files can be detected and resolved.
 Our automated data collection and analysis pipeline (Authors, 2026) supports such ongoing analyses.
-Finally, our study is limited to a single point-in-time snapshot of a rapidly evolving landscape.
-Agentic AI coding tools are under active development, and both the available configuration mechanisms and the patterns of their adoption are likely to change.
+Finally, our findings represent a point-in-time snapshot of a rapidly evolving landscape.
 The trends we identify—toward standardization around AGENTS.md, shallow adoption of advanced mechanisms, and tool-specific configuration cultures—are early empirical signals rather than settled findings.
@@
 Our study is the first to provide a comprehensive overview of the configuration mechanisms used by agentic AI coding tools in software engineering.
-We systematically identified eight configuration mechanisms across five agentic AI coding tools and analyzed their adoption in 2,926 GitHub repositories.
+We systematically identified eight configuration mechanisms across five agentic AI coding tools and analyzed their adoption in 2,923 GitHub repositories.
 Three findings stand out.
@@
 Tool providers should consider improving onboarding and documentation for advanced mechanisms, given the gap between their expressive power and current adoption.
-For researchers, controlled studies are needed to determine whether richer configuration, such as Skills with executable resources, yields measurable improvements over Context Files alone, and how conflicts between configuration artifacts in multi-tool repositories can be detected and resolved.
-Longitudinal research needs to track how configuration patterns evolve.
+For researchers, controlled studies should determine whether richer configuration yields measurable improvements over Context Files alone, and longitudinal research should track how configuration patterns evolve.

### v3

@@
 Software Engineering, Generative AI, AI Agents, Configuration
-journalyear: 2026copyright: ccdoi: XXXXXXX.XXXXXXXdoi: XXXXXXX.XXXXXXXconference: Anonymous Conference; Anoym.
-Conf.; Earthbooktitle: Anonymous Conferencedoi: 10.1145/XXXXXXX.XXXXXXXisbn: 979-8-4007-XXXX-X/2026/04ccs: Software and its engineering
+journalyear: 2026copyright: ccdoi: 10.1145/XXXXXXX.XXXXXXXconference: 3rd ACM International Conference on AI-powered Software; July 6–7, 2026; Montreal, Canadabooktitle: 3rd ACM International Conference on AI-powered Software (AIware 2026), July 6–7, 2026, Montreal, Canadaisbn: 979-8-4007-XXXX-X/2026/04ccs: Software and its engineering
 1. Introduction
@@
 A configuration artifact is a tangible instance of a mechanism (e.g., the actual Markdown file).
-For example, files like AGENTS.md or CLAUDE.md act as “READMEs for agents” with context-specific information about build commands, coding conventions, and rules for CI/CD pipelines (Mohsenimofidi et al., 2026; OpenAI, 2025).
+For example, files like AGENTS.md or CLAUDE.md act as “READMEs for agents” with context-specific information about build commands, coding conventions, and rules for CI/CD pipelines (Mohsenimofidi et al., 2025; OpenAI, 2025).
 Configuration artifacts are version-controlled, inspectable, and collaboratively maintained.
@@
 There has been no systematic examination of the broader configuration mechanisms to customize agentic behavior.
-Independent of tool architecture, context engineering allows developers to design, structure, and supply task- and project-specific information to guide the behavior of tools (Mei et al., 2025; Schmid, 2025; Horthy, 2025; Sahoo et al., 2024; DAIR.AI Prompt Engineering Guide, 2025).
+Context engineering builds on earlier prompt engineering techniques (Sahoo et al., 2024; DAIR.AI Prompt Engineering Guide, 2025) but broadens the scope from individual prompts to the systematic design of all information supplied to a tool.
+Mei et al.
+(Mei et al., 2025) survey this emerging field, while practitioner accounts demonstrate how structuring project-specific context improves agent effectiveness in complex codebases (Schmid, 2025; Horthy, 2025).
 Empirical work has mainly focused on repository-maintained context files, such as AGENTS.md.
 Mohsenimofidi et al.
-(Mohsenimofidi et al., 2026) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level context files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions.
+(Mohsenimofidi et al., 2025) analyzed the emergence of context engineering practices in open-source projects and showed that repository-level context files serve as persistent configuration mechanisms that encode architectural constraints, build commands, and workflow conventions.
 Chatlatanagulchai et al.
@@
 (Santos et al., 2025) analyzed 328 CLAUDE.md files and found that architectural concerns dominate, often co-occurring with development guidelines, testing, and project overview.
-Beyond structural characterization, recent work has begun to evaluate the impact of such artifacts.
-Lulla et al.
+Beyond structural characterization, Lulla et al.
 (Lulla et al., 2026) conducted a controlled study comparing agent executions with and without an AGENTS.md file and reported lower runtime and token consumption while maintaining comparable task completion behavior.
@@
 (Villamizar et al., 2025) argued that prompts should be treated as software engineering artifacts and outlined a research agenda for their evolution, reuse, and governance.
+Q.
 Zhang et al.
-(Zhang et al., 2025) introduced a framework that treats contexts as evolving instructions and addresses brevity bias and context collapse via incremental updates.
-Recent work further generalized this perspective by formalizing context as a programmable abstraction.
-Zhang and Wang (Zhang and Wang, 2025) introduced Monadic Context Engineering, modeling context construction and transformation through monadic composition.
+(Zhang et al., 2025a) proposed a framework treating contexts as evolving instructions, addressing brevity bias and context collapse through incremental updates.
+Other work formalizes context as a programmable abstraction.
+Y.
+Zhang et al.
+(Zhang et al., 2025b) introduced Monadic Context Engineering, modeling context construction and transformation through monadic composition.
 McMillan (McMillan, 2026) evaluated structured file-native context schemas and showed how format and multi-file organization influence retrieval accuracy and downstream agent performance.
@@
 (Ye et al., 2026) proposed Meta Context Engineering, a framework in which agentic skills and context artifacts co-evolve via evolutionary search.
-Jiang and Nam (Jiang and Nam, 2026) explored Cursor rules as one particular configuration mechanism and developed a taxonomy of project contexts.
-However, a cross-tool perspective mapping configuration mechanisms and studying their adoption patterns across repositories is missing. Our work addresses this gap.
+Jiang and Nam (Jiang and Nam, 2025) explored Cursor rules and developed a taxonomy of project contexts.
+However, no prior work maps configuration mechanisms across tools or studies their adoption patterns across repositories. We provide this cross-tool perspective.
 3. Data Collection and Analysis
-We collected OSS projects hosted on GitHub, selecting repositories belonging to “engineered” software projects (Munaiah et al., 2017) using an updated version of a selection approach from prior work, starting with the SEART GitHub search tool (Mohsenimofidi et al., 2026; SEART, 2025; Dabic et al., 2021).
+We collected OSS projects hosted on GitHub, selecting repositories belonging to “engineered” software projects (Munaiah et al., 2017) using an updated version of a selection approach from prior work, starting with the SEART GitHub search tool (Mohsenimofidi et al., 2025; SEART, 2025; Dabic et al., 2021).
 To ensure sufficient project maturity and development history, we selected non-fork repositories with at least two contributors and a license that were created before 1 January 2024 with commits since 1 June 2025.
 We then excluded archived, disabled, or locked repositories.
-We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and applied filters based on commit and watcher count based on prior research (Mohsenimofidi et al., 2026).
+We applied a licensing filter and a language filter based on popularity, selecting Python, TypeScript, JavaScript, Go, Java, C++, Rust, PHP, C#, and C, and excluded repositories with fewer than 271 commits or fewer than 7 watchers (median-based thresholds from prior research (Mohsenimofidi et al., 2025)).
 This resulted in a sample of 37,249 repositories.
@@
 (Munaiah et al., 2017), that is, showing clear evidence of software engineering practices (e.g., a stated purpose, installation and deployment instructions, documentation, testing, CI/CD) and maintenance.
-The supplementary material includes the prompt (iteratively developed and tested on subsets) and model configuration (OpenAI defaults) (Authors, 2026).
+The supplementary material includes the prompt (iteratively developed and tested on subsets) and model configuration (OpenAI defaults) (Galster et al., 2026).
 The pipeline labeled 32,564 repositories as engineered and excluded the remaining 3,620 (1,416 classified as non-engineered, 2,204 “unsure”).
@@
 This resulted in 2,923 repositories that use one or more AI coding tools (three additional repositories contained only empty configuration artifacts and were excluded).
-Our data collection and analysis scripts and the analyzed data are available online (Authors, 2026).
+Our data collection and analysis scripts and the analyzed data are available online (Galster et al., 2026).
 4. Configuration Mechanisms (RQ1)
-We selected five agentic AI coding tools (Table 2) based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026): the four most popular AI tools among the surveyed developers, substituting Codex for ChatGPT as the agentic tool by the same vendor, plus Cursor, which has recently been studied in software engineering (He et al., 2026; Jiang and Nam, 2026).
+We selected five agentic AI coding tools (Table 2) based on the 2025 Stack Overflow Developer Survey (Stack Exchange Inc., 2026): the four most popular AI tools among the surveyed developers, substituting Codex for ChatGPT as the agentic tool by the same vendor, plus Cursor, which has recently been studied in software engineering (He et al., 2025; Jiang and Nam, 2025).
 Our study focuses on the CLI-based agentic interfaces that first appeared in 2025; for some tools, prior non-agentic versions exist.
@@
 Note that for Copilot, Cursor, and Gemini, certain detected files apply to both conversational and agentic interfaces.
-All documents and the Python scripts that implement our matching strategy are part of the supplementary material (Authors, 2026).
+All documents and the Python scripts that implement our matching strategy are part of the supplementary material (Galster et al., 2026).
 {results}
@@
 Context Files (e.g., CLAUDE.md, AGENTS.md) were the most frequently adopted mechanism, used by 61.5–100% of repositories across all tools.
-Beyond Context Files, distinct tool-specific profiles emerge: 72.8% of Cursor repositories adopt Rules, since Cursor was one of the first to introduce this mechanism (Jiang and Nam, 2026), and 61.2% of Gemini repositories use Settings.
+Beyond Context Files, distinct tool-specific profiles emerge: 72.8% of Cursor repositories adopt Rules, since Cursor was one of the first to introduce this mechanism (Jiang and Nam, 2025), and 61.2% of Gemini repositories use Settings.
 Other mechanisms remain below 20% of repositories for most tools.
@@
 Context Files are Markdown files that provide central, machine-readable context about a project.
-AGENTS.md, introduced by OpenAI, now serves as an open tool-agnostic convention increasingly supported by various tools (Mohsenimofidi et al., 2026).
+AGENTS.md, introduced by OpenAI, now serves as an open tool-agnostic convention increasingly supported by various tools (Mohsenimofidi et al., 2025).
 We identified 4,854 Context Files across 2,631 of the 2,923 repositories in our sample.
@@
 Internal validity:
-We classified repositories as “engineered” software projects using GPT-5.2 on their README content, with iterative prompt development and spot-checks to mitigate errors.
-We used a single labeling run without considering inter-model agreement; 2,204 “unsure” cases (often due to inaccessible linked resources) were excluded.
+We classified repositories as “engineered” software projects based on their README content.
+During iterative prompt development, we tested GPT-5.2, GPT-5-nano, and GPT-5-mini, and selected GPT-5.2 because spot-checks showed it produced fewer “unsure” labels and fewer false positives and negatives.
+We used a single labeling run; 2,204 “unsure” cases (often due to inaccessible linked resources) were excluded.
 Future work should assess the reliability of this approach by adding alternative models and configurations.
@@
 We selected repositories showing established engineering practices, but cannot claim representativeness for closed-source development, nor did we examine variation across application domains.
-Finally, our findings are a point-in-time snapshot (February 2026) of a rapidly evolving landscape.
+Finally, our findings are a point-in-time snapshot (February 2026) of a fast-moving field.
 7. Discussion
-Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,923 repositories.
-Below, we discuss the implications of our study.
+Our study offers a first cross-tool snapshot of configuration mechanisms and their artifacts across five agentic AI coding tools and 2,923 GitHub repositories.
+Below, we discuss the implications of our findings.
 Standardization around AGENTS.md:
@@
 Configuration practices vary systematically across tools: Claude Code repositories exhibit the broadest configuration footprint, Cursor projects emphasize Rules and Commands, and Copilot repositories rarely extend beyond Context Files.
-These differences likely reflect both the configuration options each tool exposes (Table 1) and the norms emerging within their respective user communities.
+These differences likely reflect the configuration options each tool exposes (Table 1) and the norms emerging within their user communities.
 Repository characteristics also differ: Cursor repositories tend to be younger and larger, while Gemini repositories show higher activity levels (Table 3).
@@
 Finally, adopting tool-specific mechanisms (e.g., rule-based configurations) may shape workflows in tool-dependent ways.
-For researchers, our findings point to several directions.
-Longitudinal studies are needed to track how configuration practices evolve as tools mature.
-Controlled studies should assess whether advanced mechanisms such as Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone (Lulla et al., 2026).
-Given that certain projects use multiple tools (Section 5.1), research should also investigate interoperability challenges, including how conflicts between overlapping or contradictory instructions across files can be detected and resolved.
-Our automated data collection and analysis pipeline (Authors, 2026) supports such ongoing analyses.
-Finally, our findings represent a point-in-time snapshot of a rapidly evolving landscape.
+For researchers, our findings motivate longitudinal studies to track how configuration practices evolve as tools mature, and controlled studies to assess whether Skills with executable resources or dedicated Subagents provide measurable benefits over Context Files alone (Lulla et al., 2026).
+Given that certain projects use multiple tools (Section 5.1), research should also investigate how conflicts between overlapping instructions across files can be detected and resolved.
+Our data collection and analysis pipeline (Galster et al., 2026) supports such ongoing analyses.
+Finally, our findings represent a point-in-time snapshot of a fast-moving field.
 The trends we identify—toward standardization around AGENTS.md, shallow adoption of advanced mechanisms, and tool-specific configuration cultures—are early empirical signals rather than settled findings.
 8. Conclusions
-Our study is the first to provide a comprehensive overview of the configuration mechanisms used by agentic AI coding tools in software engineering.
-We systematically identified eight configuration mechanisms across five agentic AI coding tools and analyzed their adoption in 2,923 GitHub repositories.
+Our study provides the first comprehensive overview of configuration mechanisms for agentic AI coding tools, identifying eight mechanisms across five tools and analyzing their adoption in 2,923 GitHub repositories.
 Three findings stand out.
@@
 Third, distinct configuration cultures are forming around different tools.
-For practitioners, these findings suggest that AGENTS.md is the natural starting point for configuring agentic tools, especially in multi-tool environments, and that Skills offer untapped potential for encoding recurring workflows beyond static instructions.
-Tool providers should consider improving onboarding and documentation for advanced mechanisms, given the gap between their expressive power and current adoption.
-For researchers, controlled studies should determine whether richer configuration yields measurable improvements over Context Files alone, and longitudinal research should track how configuration patterns evolve.
+For practitioners, AGENTS.md is the natural starting point for configuring agentic tools, especially in multi-tool environments, and Skills offer untapped potential for encoding recurring workflows beyond static instructions.
+Tool providers should improve onboarding for Skills and Subagents, which few repositories use despite their ability to encapsulate executable workflows and isolated agent contexts.
+For researchers, controlled studies should determine whether richer configuration improves outcomes over Context Files alone, and longitudinal research should track how these patterns evolve.